### v4

# A variational formulation of the free energy of mixed quantum-classical systems: coupling classical and electronic density functional theories

Abstract
Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures. However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations.
This paper aims to develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation and serves as a starting point for further development. We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble. We start with the adiabatic equilibrium density matrix for a system comprising $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles and, based on this, we propose a variational formulation of the Helmholtz free energy in terms of the full QM/MM density matrix. Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities, thus generalizing both cDFT and electronic DFT (eDFT) for QM/MM systems.
We then reformulate the functional in order to explicitly make appear the eDFT and cDFT Levy-Lieb functionals, together with a new universal correlation functional for QM/MM systems. A mean-field approximation is finally introduced in the context of solvation problems and we discuss its connection with several existing mixed cDFT-eDFT schemes.
I Introduction
The Hohenberg-Kohn theoremsHohenberg and Kohn (1964) dramatically reduce the complexity of simulating a quantum (QM) molecular system by enabling a focus on the one-electron density instead of the full $N$-electron wave function. Central to this simplification is the unknown universal density functional, which was later given a more rigorous mathematical foundation by Levy and LiebLevy (1979); Lieb (1983). Through the self-consistent framework introduced by Kohn and ShamKohn and Sham (1965), along with extensive efforts to develop approximations to the exact exchange-correlation functional (see, for example, a recent review in Ref. Toulouse, 2022), electronic density functional theory (eDFT) has become the workhorse of quantum chemistry.
Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications. Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion. In this context, it is natural to compute the forces acting on the nuclei at the quantum level and then to solve the classical equations of motion to propagate the trajectory. This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999). Despite its considerable achievements, it is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space. The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al. (2013); Rozsa et al. (2018). To alleviate ab-initio molecular dynamics computational burden, QM/MM methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM). The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion. However, even in QM/MM approaches, sampling the phase space of large systems remains computationally intensive.
Inspired by the extension of eDFT to QM systems at finite temperaturesMermin (1965), density-based formalisms have emerged for describing purely classical objects in statistical mechanics, particularly in the form of classical density functional theory (cDFT)Evans (1979). Originally rooted in liquid state theory, cDFT is essentially a classical counterpart to eDFT, reformulating the classical statistical mechanics problem of large ensembles of identical particles (atoms or molecules) in terms of one-body density functionals. cDFT, which is naturally formulated within the grand canonical ensemble, establishes the existence of a unique functional of the classical particle density that reaches its minimum at the equilibrium density. At this minimum, the functional is equal to the grand potential, which serves as the thermodynamic potential of the grand canonical ensemble. Therefore, most equilibrium properties can be obtained through functional minimization over the classical one-body density rather than through tedious sampling of configurations in an exponentially large phase space.
The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object—such as the wavefunction in QM or the probability distribution in MM—with the optimization of a functional that depends only on a low-dimensional quantity, i.e. the one-body density of either quantum electrons or classical molecules. The mathematical framework allowing for such a simplification is the so-called constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFT, although not usable in practical calculations.
Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al. (2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024). In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians.However, the QM/MM interaction lacks of a unique practical definition. The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023). The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals. While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007) justified their so-called joint density-functional formulation by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei. To derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei. Howvever, this step makes it difficult to pursue a rigorous derivation and their working functional is rather defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals. In other workJeanmairet, Levesque, and Borgis (2020); Tang et al. (2020), a similar form of the functional was invoked without further theoretical justification.
Thus, while QM/cDFT approaches offer potentially attractive computational scaling, their derivation from an exact theory remains somewhat ambiguous. This work addresses this gap by establishing a rigorous mathematical framework that provides a DFT-equivalent description of mixed QM/MM systems. Here, we focus on the canonical ensemble, deriving equations to compute the Helmholtz free energy using only the QM and MM one-body densities. Extending this formulation to the grand canonical ensemble for classical particles is conceptually straightforward but introduces additional complexity in the notations due to the varying number of particles. The main difference between the two ensembles is that, in the grand canonical ensemble, there exists a unique correspondence between the equilibrium density and the external potential. In contrast, in the canonical ensemble, the equilibrium density corresponds to a family of external potentials, defined up to an additive constantLutsko (2022).
Within the present paper, the foundational object from which the QM/MM DFT formulation is derived is the explicit expression of the adiabatic equilibrium density of a mixed QM/MM system in the canonical ensemble. This expression, which was previously established in the adiabatic basis by Nielsen et. al. in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), is recovered here following an alternative route (see Sec. II of the supplementary materials). Building on this, we propose a functional formulation of the free energy in terms of general QM and MM high-dimensional densities (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for a proof of the variational properties). By applying the Levy-Lieb constrained search and assuming additive two-body QM/MM interactions, we then naturally express the free energy in terms of the one-electron density and the classical one-body density. We explicitly establish the connection between eDFT and cDFT through the introduction of a new universal QM/MM correlation one-body density functional. Additionally, we discuss a mean-field approximation and its relationship with previously proposed cDFT/eDFT couplings in the context of QM/MM solvation problems. We also note that this work can be seen as the QM/MM and finite-temperature counterpart of the recently published study by Fromager and LasorneFromager and Lasorne (2024), which formulated an exact DFT framework incorporating both electrons and nuclei at the quantum level—going beyond the Born-Oppenheimer approximation.
The paper is organized as follows. In Sec. II, we begin by discussing the adiabatic equilibrium density for a mixed QM/MM system in the canonical ensemble (Sec. II.1 and II.2). Then, in Sec. II.3, using the explicit form of the equilibrium density, we reformulate the computation of the free energy as a functional minimization over general high-dimensional mixed QM/MM densities. Sec. III focuses on expressing this problem explicitly in terms of the one-body QM and MM densities. In Sec. III.1, we employ the Levy-Lieb constrained search to reformulate the high-dimensional QM/MM functional problem by recognizing a many-to-one correspondence between the high-dimensional QM/MM quantities and the one-body QM and MM densities. In Sec. III.2, we establish a connection between eDFT and cDFT by introducing a universal QM/MM correlation functional of the one-body QM and MM densities. Sec. III.3 then discusses this formulation in the context of solvation problems and recent QM/cDFT coupling approaches. Finally, Sec. IV summarizes the main results and conclusions. The supplementary materials provide an alternative derivation of the equilibrium density for a QM/MM system that does not rely on Liouvillian operators, and the appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ provides a proof of the variational properties of the QM/MM functional.
II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
The main goal of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables. We define in Sec. II.1 our working Hamiltonian describing the QM/MM system. The equilibrium density of a mixed QM/MM system is then given in Sec. II.2, and we briefly discuss the physical meaning of that result. We eventually show in Sec. II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
II.1 The Hamiltonian of the quantum-classical system
We adopt the following notation conventions : i) lowercase letters will refer to purely quantum objects (both variables and operators) while capital letters will refer to either purely classical or mixed quantum-classical objects and ii) all bold letters refer to collective variables or operators (i.e. relative to all quantum or classical particles) and non bold letters will refer to individual or pairwise variables or operators.
The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$, and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
The Hamiltonian of the total system is written as
$\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P},{\bf{q}})=\hat{\bf h}_{{\text{qm}}}({\bf{q}})+{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}}).$ (1)
In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous. In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons) in interaction with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$ local in position representation, and interacting among themselves with a pairwise potential ${w}_{{\text{qm}}}({{q}}_{1},{{q}}_{2})$. The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
$\hat{\bf h}_{{\text{qm}}}=\hat{\bf t}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+\hat{\bf w}_{{\text{qm}}},$ (2)
with
$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\hbar\nabla_{{{q}}_{i}},$ (3)
$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{j}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
Then, ${\bf H}_{{\text{mm}}}$ is the classical Hamiltonian function representing $N_{{\text{mm}}}$ identical classical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves with a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$. These particles are typically effective particles such as atoms or rigid molecules instead of the individual electrons composing them and they interact through model, parametrized, force fields. The explicit form of the classical Hamiltonian ${\bf H}_{{\text{mm}}}$ reads then
${\bf H}_{{\text{mm}}}={\bf T}+{\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (5)
with
${\bf T}=\sum_{i=1}^{N_{{\text{mm}}}}\frac{{{P}}_{i}^{2}}{2M},\qquad{\bf V}^{\text{ext}}_{{\text{mm}}}=\sum_{i=1}^{N_{{\text{mm}}}}V^{\text{ext}}_{{\text{mm}}}({{Q}}_{i}),$ (6)
${\bf W}_{{\text{mm}}}=\sum_{j=1}^{N_{{\text{mm}}}}\sum_{i>j}{W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j}).$ (7)
Regarding the coupling term ${W}_{{\text{qm}}}^{{\text{mm}}}$, we simply assume here that it is an additive pairwise potential, i.e.
$\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}=\sum_{i=1}^{N_{{\text{qm}}}}\sum_{j=1}^{N_{{\text{mm}}}}{W}_{{\text{qm}}}^{{\text{mm}}}({{q}}_{i},{{Q}}_{j})$ (8)
The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position ${Q}$ and a fundamental particle (such as an electron) at position ${q}$.
II.2 Mixed quantum-classical equilibrium density
The equilibrium density corresponding to the Hamiltonian of Eq. (1) can be written as
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}},$ (9)
where $\beta=(\text{k}_{\text{B}}T)^{-1}$, $\text{k}_{\text{B}}$ is the Boltzmann constant, and the dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables ${\bf{Q}}$ is here explicitly highlighted. This result is nothing but the real-space representation of that obtained by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis. The result of Eq. (9) can be obtained by starting from an Hamiltonian where both sets of particles are considered to be quantum. Then, one performs a Wigner transform over the $N_{{\text{mm}}}$ particles only and eventually a semi-classical limit is obtained by letting the ratio of masses $(m/M)^{1/2}$ tends towards 0. Conserving only the zeroth order term of this semi-classical limit yields to Eq. (9). This expansion in terms of the ratio of the masses is natural since the quantum particles, typically electrons, are usually much lighter than the classical ones, typically atoms or molecules. The original derivation of Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001) used the Liouvillian imaginary time evolution equation of the density matrix together with the Poisson-brakets operatorsİmre et al. (1967), which is a well-known formalism in the quantum dynamics communityKapral and Ciccotti (1999); Nielsen, Kapral, and Ciccotti (2001); Kapral (2006, 2015). However, we provide in the supplementary materials an alternative derivation which involves a similarity transformation and resorts to the Baker-Campbell-Hausdorf expansion to express the Wigner transform Hamiltonian as a series of nested commutators, which might appear more familiar to the ”static” quantum chemistry community.
In Eq. (9), the normalization is the partition function of the QM/MM system which is expressed as
$Z^{\text{tot}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}}\int\text{d}{\bf{Q}}\text{d}{\bf P}\!\!\int\!\text{d}{\bf{q}}$ ${\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}$ (10)
$e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}.$
The corresponding Helmholtz free energy is given by
$F_{0}^{\text{tot}}=-\text{k}_{\text{B}}T\log(Z^{\text{tot}}).$ (11)
While the expressions of Eqs. (9) and (10) are given in real-space, they can easily be written in the adiabatic basis, where one explicitly sees that the density matrix is diagonal as found by Nielsen et. al.Nielsen, Kapral, and Ciccotti (2001).
Although the expression of the equilibrium density in Eq. (9) is relatively simple, it calls some comments. Firstly, as usual in classical statistical mechanics, the contribution to the Helmholtz free energy coming from the purely classical kinetic term in ${\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})$ can be computed analytically. Therefore it is worth splitting the equilibrium density in terms of ${\bf{Q}}$ and ${\bf P}$ dependence
$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{e^{-\beta{\bf T}}}{z_{{\bf P}}}\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}),$ (12)
where ${\bf T}$ is the purely classical kinetic energy defined in Eq. (6), and
$\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}}{Z},$ (13)
${{\bf{V}}}_{\text{mm}}$ $={\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (14)
$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg{(}\frac{2\pi M}{\beta h^{2}}\bigg{)}^{\frac{3}{2}N_{{\text{mm}}}},$ (15)
$Z$ $=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})},$ (16)
$Z^{\text{tot}}$ $=z_{{\bf P}}Z,$ (17)
such that the free energy can be written as
$F_{0}^{\text{tot}}$ $=F_{0}+F_{{\bf P}},$ (18)
where
$F_{{\bf P}}=-\text{k}_{\text{B}}T\log(z_{{\bf P}}),$ (19)
and
$F_{0}=-\text{k}_{\text{B}}T\log(Z)$ (20)
is the interacting part of the Helmholtz free energy of the QM/MM system.
Secondly, the object $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ in Eq.(13) can be expressed as the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
$\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})={\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})\boldsymbol{p}_{\text{eq}}({\bf{Q}}),$ (21)
with
${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ $={\langle{\bf{q}}|}\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}}){|{\bf{q}}^{\prime}\rangle},$ (22)
$\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})=\frac{1}{Z^{{\text{qm}}}({\bf{Q}})}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))},$ (23)
$Z^{{\text{qm}}}({\bf{Q}})$ $=\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle},$ (24)
$\boldsymbol{p}_{\text{eq}}({\bf{Q}})=\frac{Z^{{\text{qm}}}({\bf{Q}})}{Z}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (25)
$Z$ and $Z^{{\text{qm}}}({\bf{Q}})$ are related through the following equation
$Z$ $=\int\text{d}{\bf{Q}}Z^{{\text{qm}}}({\bf{Q}})e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (26)
It is noteworthy that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ defined in Eq. (22),(23) is a positive definite operator and that $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$ of Eq. (25) is also positive definite and normalized to unity. Therefore, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are simply a special case of more general density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively. This will allow us in Sec. II.3 to propose a variational formulation of the Helmholtz free energy for a QM/MM system.
Additionally, this decomposition of the equilibrium density as a product of a purely quantum equilibrium density and a purely classical equilibrium density could lead to the somehow strange conclusion that there is no correlation between the quantum and classical systems, even if the two systems interact. Nevertheless, one must pay attention to the fact that the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$. Therefore, the two systems are indeed correlated.
Thirdly, an obvious illustration of the meaning of the equilibrium density of Eq. (13) is that corresponding to the Born-Oppenheimer approximation. In the latter case, the light particles described by ${\bf{q}}$ are electrons and the heavy particles described by ${\bf{Q}}$ are the classical nuclei. There, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, there is no external potential both for the QM and MM parts, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is the nuclei-electron attraction, and ${\bf W}_{{\text{mm}}}({\bf{Q}})$ is the Coulomb repulsion between the nuclei, such that the operator happening in the exponential of Eq. (13) is the Born-Oppenheimer Hamiltonian.
A follow up on this comment is that, as mentioned earlier, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is diagonal in the adiabatic basis, which is equivalent to say that non-adiabatic couplings are completely neglected. In the context of ab initio MD, this implies that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ can be obtained by averaging, over all possible adiabatic electronic states, the classical trajectories of nuclei computed from the forces of a given adiabatic electronic state. Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is referred to as the adiabatic QM/MM density matrix. There is a flourishing literature concerned with going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM treatments (see for instance Ref. Kapral, 2015 for a review), but the treatment of non adiabatic couplings goes beyond our QM/MM scheme.
II.3 Reformulation of the Helmholtz free energy for a QM/MM system within a variational framework
II.3.1 Variational formulation
As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we want to reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq. (21) is the minimizer of that functional and that it yields the correct Helmholtz free energy of Eq. (20). Moreover, since the ultimate goal is to formulate a variational principle involving both quantum and classical densities, it is natural to use the decomposition of Eq. (21).
Thus, we would like to write the Helmholtz free energy as the minimum of a functional defined on a couple of quantum and classical objects as follows
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (27)
where the operator $\hat{{\boldsymbol{\rho}}}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
$\hat{{\boldsymbol{\rho}}}\equiv\hat{{\boldsymbol{\rho}}}({\bf{Q}}),$ (28)
and $\boldsymbol{p}$ is a classical probability density representing $N_{{\text{mm}}}$ classical particles and not depending on any quantum variable. Therefore, ${\boldsymbol{\rho}}$ and $\boldsymbol{p}$ must be normalized to be density matrices and probability measures, respectively, i.e
$\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})=1\quad\forall\,\,{\bf{Q}},$ $\int\text{d}{\bf{Q}}\,\,\boldsymbol{p}({\bf{Q}})=1,$ (29)
where ${\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ is the real-space representation of the operator $\hat{{\boldsymbol{\rho}}}$, i.e.
${\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})={\langle{\bf{q}}|}\hat{{\boldsymbol{\rho}}}({\bf{Q}}){|{\bf{q}}^{\prime}\rangle},$ (30)
such that that the product of the two functions is normalized to unity
$\text{Tr}\{\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\}=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (31)
Therefore, the equilibrium density of the mixed QM/MM system described in Eq. (21) lies within the functional space defined by all the couple of functions and density matrices satisfying Eqs. (29)-(31). As shown in the appendix (see section Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$), if we define the Helmholtz functional as follows
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (32)
where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has was accounted for in $F_{{\bf P}}$)
$\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}(\hat{\bf h}_{{\text{qm}}}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+{\bf W}_{{\text{mm}}}+{\bf V}^{\text{ext}}_{{\text{mm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}})\big{\}},$ (33)
and $\mathcal{S}$ is the dimensionless entropy of the total system
$\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=-\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\log\big{(}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\big{)}\big{\}},$ (34)
then the only minimizer of the functional of Eq. (32) is $(\hat{{\boldsymbol{\rho}}}_{\text{eq}},\boldsymbol{p}_{\text{eq}})$ of Eq. (21), i.e
$-\text{k}_{\text{B}}T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}},\boldsymbol{p}),$ (35)
and
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=F_{0}=-\text{k}_{\text{B}}T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (36)
II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$
Because $\boldsymbol{p}$ does not depend on the quantum variables and that we chose the normalization condition of Eq. (29), we can naturally decompose the entropy as
$\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}],$ (37)
with $S_{{\text{mm}}}[\boldsymbol{p}]$ being the purely classical dimension-less entropy
$S_{{\text{mm}}}[\boldsymbol{p}]$ $=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\log(\boldsymbol{p}({\bf{Q}})),$ (38)
and $S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the quantum-classical dimension-less entropy
$S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=-\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\log(\hat{{\boldsymbol{\rho}}})\big{\}}$ (39)
$=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\log({\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})),$
which depends explicitly on both $\hat{{\boldsymbol{\rho}}}$ and $\boldsymbol{p}$.
We can similarly decompose $\mathcal{U}$ as follows
$\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ (40)
$+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$
where $E_{\text{mm}}[\boldsymbol{p}]$ is the mutual interaction of the classical part
$E_{\text{mm}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf W}_{{\text{mm}}}({\bf{Q}}),$ (41)
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ is the functional accounting for the interaction of the classical particles with the external potential
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}}),$ (42)
$E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the kinetic and mutual interaction of the quantum part
$E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=\text{Tr}\{(\hat{\bf t}+\hat{\bf w}_{{\text{qm}}})\boldsymbol{p}\hat{{\boldsymbol{\rho}}}\}$ (43)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})$
$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+{w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the functional accounting for the interaction of the quantum particles with the external potential
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=\text{Tr}\{\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}\boldsymbol{p}\hat{{\boldsymbol{\rho}}}\}$ (44)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}{\bf v}^{\text{ext}}_{{\text{qm}}}({\bf{q}},{\bf{q}}){\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}}),$
and eventually $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is the quantum-classical interaction,
$E_{{\text{qm}}}^{{\text{mm}}}$ $[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\text{Tr}\{\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}\boldsymbol{p}\hat{{\boldsymbol{\rho}}}\}$ (45)
$=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\int\text{d}{\bf{q}}^{\prime}{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}){\boldsymbol{\rho}}({\bf{q}}^{\prime},{\bf{q}},{\bf{Q}}).$
It is noteworthy here to highlight that the form of Eq. (45) is quite general as it consists in a priori in potential with a non local kernel ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$. This form of interaction includes for instance Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
Eventually, once all terms are inserted in Eq. (27), the Helmholtz free energy $F_{0}$ can be obtained as
[…]
$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\big{(}S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big{)}\big{\}}.$
Eq. (46) is therefore a variational formulation of $F_{0}$ for a mixed QM/MM system in terms of a minimization over the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$, which, to the best of our knowledge, was never formalized before. One can see from Eqs. (38), (41) and (42) that the functionals corresponding to the MM part of the energy and entropy are completely independent of the quantum part of the system, as they depend only on the classical probability $\boldsymbol{p}$. Nevertheless, because of the term $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$, the quantum and classical parts are coupled in the minimization process of Eq. (46). Also, one can see that the functionals corresponding to the QM parts of the energy and entropy explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
Although Eq. (46) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the variables $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ involved in the minimization are nevertheless very high-dimensional. Therefore, the goal of the following sections is to reformulate the variational principle of Eq. (46) using functionals of significantly lower-dimensional densities, i.e. the one-body QM and MM densities.
III One-body density formulation through Levy-Lieb constrained search formulation
An important consequence of the one-body nature of both $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, along with the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}}$ is that the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ in Eqs. (42) and (44) can be computed using one-body objects. Therefore, by applying the Levy-Lieb constrained search, we can reformulate the variational problem of Eq. (46) in terms of a functional that depends only on the one-body quantum and classical densities, rather than the full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables.
III.1 Classical and quantum one-body density and the Levy-Lieb constrained search
We naturally introduce the one-body classical density $n({Q})$ as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq. (29), is simply
[…]
Similarly we can define the QM one-body density $\rho({q})$ as
[…]
It is noteworthy from Eq. (48) that, because of the ${\bf{Q}}$ dependence in $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$, the classical $N$-body probability $\boldsymbol{p}({\bf{Q}})$ contributes to the one-body quantum density $\rho({q})$, whereas from Eq. (47) it is clear that the quantum $N$-body density matrix $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ does not contribute to the classical one-body density.
By using permutation symmetry of both $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to exchange of either QM or MM variables, we can then write the two terms arising from the external potential as functionals of the QM and MM one-body densities
$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}})$ (49)
$=\int\text{d}{Q}n({Q})V^{\text{ext}}_{{\text{mm}}}({Q})$
$\equiv(V^{\text{ext}}_{{\text{mm}}}|n),$
$E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}{\bf v}^{\text{ext}}_{{\text{qm}}}({\bf{q}},{\bf{q}}){\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})$ (50)
$=\int\text{d}{q}\rho({q})v^{\text{ext}}_{{\text{qm}}}({q})$
$\equiv(v^{\text{ext}}_{{\text{qm}}}|\rho).$
As the contributions coming from the external potentials can be computed directly with a couple of one-body object $(\rho,n)$, we will now introduce a universal functional of $(\rho,n)$ which contains all necessary information to compute the Gibbs free energy. This functional is built using the Levy-Lieb constrained search formalism as in eDFT. First, we introduce the intrinsic Gibbs free energy functional of the $N$-body objects
$f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}].$ (51)
such that we can then rewrite the Gibbs free energy as
$F_{0}=\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}\big{(}E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]+f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\big{)}.$ (52)
Then, the many-to-one correspondence between $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ and $(\rho,n)$ allows us to introduce the Levy-Lieb constrained search as follows
$\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\min_{(\rho,n)}\,\,\min_{(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (53)
where the notation $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is done over the set of couple of $N$-body objects $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ having the couple of 1-body object given by $(\rho,n)$. Therefore, by remembering that the external parts depends only on the one-body densities (see Eqs. (49) and (50)), one can rewrite Eq. (53) as
$\min_{\hat{{\boldsymbol{\rho}}},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=$ $\min_{(\rho,n)}\big{[}(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)$ (54)
$+$ $\min_{(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\big{]}.$
Therefore, we can introduce the universal Levy-Lieb functional of the one-body classical and quantum density as
$\mathcal{F}[\rho,n]=\min_{(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (55)
such that one can rewrite the Gibbs free energy as
$F_{0}=\min_{(\rho,n)}\big{(}(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big{)},$ (56)
which achieves to prove that we can obtain $F_{0}$ as a minimization over only one-body densities.
III.2 Explicit coupling in terms of eDFT and cDFT
Although Eq. (56) allows us to obtain the Gibbs free energy only in terms of the one-body densities $(\rho,n)$, such a functional remains unknown and the explicit connection with well known universal quantum and classical density functionals is yet unclear. This paragraph is concerned with a rewriting of $\mathcal{F}[\rho,n]$ such that the functional of MerminMermin (1965) and EvansEvans (1979) explicitly appear, together with a new QM/MM correlation functional.
We begin by recalling the universal purely quantum one-body density functionals of Mermin
$\mathcal{F}_{\text{qm}}[\rho]=\min_{\hat{{\boldsymbol{\rho}}}\rightarrow\rho}\text{Tr}\{\hat{{\boldsymbol{\rho}}}\big{(}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}})\big{)}\},$ (57)
where $\hat{{\boldsymbol{\rho}}}$ is a quantum $N$-body density matrix and $(\hat{{\boldsymbol{\rho}}}\rightarrow\rho)$ means that the minimization is performed over $N$-body quantum density matrices $\hat{{\boldsymbol{\rho}}}$ leading to a given quantum one-body density $\rho$. Similarly, we can define the universal classical one-body functional of Evans
$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big{(}{\bf W}_{{\text{mm}}}+\text{k}_{\text{B}}T\log(\boldsymbol{p})\big{)}\},$ (58)
where $\boldsymbol{p}$ is a classical $N$-body probability distribution and $(\boldsymbol{p}\rightarrow n)$ means that the minimization is performed over $N$-body classical probability distribution $\boldsymbol{p}$ leading to a given classical one-body density $n$. Historically, both Mermin and Evans introduced their universal functionals in the Grand Canonical ensemble, so Eqs. (57) and (58) are actually the Canonical counter part of the original functionals. Also, in their respective original papersMermin (1965); Evans (1979), neither Mermin or Evans explicitly used the constrained search framework which was formalized later by Levy and LiebLevy (1979); Lieb (1983).
We can then rewrite $\mathcal{F}[\rho,n]$ as follows
$\mathcal{F}[\rho,n]=$ $\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n],$ (59)
where $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the mean-field QM/MM interaction
$\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ $=\int\text{d}{q}\int\text{d}{Q}{W}_{{\text{qm}}}^{{\text{mm}}}({Q},{q})n({Q})\rho({q}),$ (60)
which will be the QM/MM equivalent of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional defined of course as the complementary functional
$\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]=\mathcal{F}[\rho,n]-\big{(}\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{)}.$ (61)
The functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq. (61) is there to reproduce all correlation terms arising because of the QM/MM interaction. The latter contains therefore some energetic and entropic terms.
We can then rewrite explicitly $F_{0}$ as
[…]
$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{\}},$
where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms. If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
III.3 Applications to solvation problems
The development presented here were motivated by the necessity of a well grounded theoretical framework for a DFT description of a quantum solute solvated in a classical solvent. This section discusses the application of Eq. (62) to this particular problem.
As common in electronic DFT calculations with an implicit solvent description, the nuclei of the molecular solute are considered clamped. The Hamiltonian of the quantum system of Eq. (2) is therefore the Born-Oppenheimer Hamiltonian.
Regarding now the classical Hamiltonian, it strongly depends on the nature of the considered solvent. We simply assume here that the external part of this potential, ${\bf V}^{\text{ext}}_{{\text{mm}}}$, represents the interaction between the solvent and the classical nuclei of the solute, and that it is local and expressed as a sum of one-body terms. Therefore, ${\bf V}^{\text{ext}}_{{\text{mm}}}$ naturally contains a Coulombic term but additional terms can be incorporated, in order, for instance, to model repulsion and dispersion interactions. There exists multiple types of classical potential modeling dispersion and repulsion, the most popular one being the Lennard-Jones potential.
Central to the description of the QM/MM interaction, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$, accounts for the electron-solvent interaction. This is a part of the total Hamiltonian that is both unknown and whose quality is critical since it essentially accounts for the information that has been lost when replacing quantum particles by classical particles. This interaction is often modeled using a Coulomb potential. Since only the external potentials are required to be local to derive Eq. (62), more sophisticated models incorporating non-local quantum operators can also be incorporated in $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$. This is a route that should be explored to improve the quality of the QM/MM model.
We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $\text{k}_{\text{B}}T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq. (57)). This is equivalent to considering only the ground state of the electronic quantum system. To illustrate the validity of this approximation for systems at room temperature ($T=300\,K$), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$. For a gap of 1.63 eV ($\approx 0.06\,\text{Ha}$), which marks the onset of the visible spectrum, this ratio drops dramatically to $10^{-28}$. Since most chemical systems near their equilibrium geometry exhibit electronic gaps of several eV, this approximation is well-suited for typical room-temperature conditions. As a consequence, we can take the limit $T\rightarrow 0$ in the definition of the Mermin universal functional (see Eq. (57)) to recover the usual universal Levy-Lieb electronic density functional, namely
$\mathcal{F}_{\text{qm}}[\rho]\approx\min_{\Psi\rightarrow\rho}{\langle\Psi|}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}{|\Psi\rangle}.$ (63)
Regarding solvent effects, we focus on cases where the solvent-solute interaction is not too strong, meaning there is no covalent bonding between the solute and solvent species. Under these conditions, the previous $T=0~{}K$ assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
In previous work about mixed electronic-classical density functional theory to study QM/MM systems, the classical fluid was described in the Grand Canonical ensemble. This is because classical DFT is naturally formulated in this ensemble. In practice this means that the number of classical particle is now allowed to fluctuate by exchanging with a reservoir of fixed chemical potential $\mu$. In this ensemble, Eq. (62)
[…]
$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{\}}$
where $\Omega_{0}$ is the grand potential, that is the state function in the grand canonical ensemble, and $\mathcal{F}_{\text{mm}}$ is now the grand canonical universal functional a la Evans.
In Eq. (64) one recognizes the usual quantum Born-Oppenheimer part for the electron (i.e. $(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$), together with the purely classical part of the solvent (i.e. $\mathcal{F}_{\text{mm}}[n]+(v_{nucl}^{{\text{mm}}}-\mu|n)$), and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for the mean-field electron-solvent interaction, while $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for solvation correlation effects in the QMMM system.
In practice, Eq. (64) requires some approximations. First, the choice of the functionals describing the quantum system and the liquid considerably influences the quality of the predictions. However, determining the appropriate electronic functional to describe a given solute and the suitable functional for a classical fluid are topics that themselves deserve a thorough review. Therefore, we let this discussion aside to focus on the QM/MM coupling. In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential. Thus Eq. (64) solely contains the quantum functional $\mathcal{F}_{\text{qm}}$, the solvent one $\mathcal{F}_{\text{mm}}$ and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$. The coupling term is modeled as a dielectric screening due to a non local dielectric function depending on the electron density and a mean-field term between both densities interacting through a Gaussian repulsive potential.
In Tang et alTang et al. (2020) Reaction DFT, both electronic and classical densities are not optimized simultaneously. Instead, the quantum functional is optimized in a continuum model. The ground state electronic density is used to compute point charges on the solute atoms which generate an external electrostatic potential in the subsequent classical functional optimization. The interaction between the quantum and the classical parts is similar to Petrosyan’s work, i.e the contribution of the electron and nuclei are gathered together. It is limited to the mean-field term, here a Coulombic term between point charges and a Lennard-Jones term and the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ is lacking. Moreover, since both density are not optimized self-consistently the interaction term $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$ only influences the classical optimization and not the quantum one.
In our previous worksJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq. (62) and limit ourselves to the mean-field term. The interaction potential was also the sum of a Lennard-Jones and a Coulombic term. However, we do not resort to point charges representation to compute the electrostatic energy but directly use the electronic density
[…]
where $n_{C}$ is the solvent charge density.
We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.62.
The results obtained in these worksPetrosyan, Rigos, and Arias (2005); Petrosyan et al. (2007); Tang et al. (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
IV Conclusion
In the present work we have developed exact variational formulations of the Helmholtz free energy of a QM/MM system made of $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ identical particles. An important result is that the exact Helmholtz free energy of a QM/MM system can be obtained though a minimization of functionals over objects of reduced dimensions, namely the one-body classical and quantum density.
The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec. II.3). This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec. Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for proofs of the variational property). The latter was proposed thanks to the knowledge of the equilibrium density of the QM/MM system (see Eqs. (13) and (16)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref. Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
Starting from this QM/MM variational principle written in terms of $N$-body objects and thanks to constrained search formalisms similar to that proposed by Levy and Lieb in eDFT, we establish a universal QM/MM functional depending only on the one-body classical and quantum objects (see Sec. III.1). Building on these results, we explicitly make a coupling with the finite temperature eDFT and cDFT functionals, which allows us to rigorously link the two theories for the description of a QM/MM system (see Sec. III.2). We then propose a practical application of this theory in the case of solvation problems, and propose a mean-field treatment. Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
Supplementary materials
The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble and II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators.
Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$
We wish to show that i) the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ of Eq. (32) is bounded from below by the Helmholtz free energy $F_{0}=-\text{k}_{\text{B}}T\log(Z)$ with $Z$ defined in Eq. (16), and ii) that the only minimizer of $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ is $\mathcal{P}_{\text{eq}}$ of in Eq. (13). We will use Jensen’s inequality from convex analysis and probability theory to prove these two statements straightforwardly. This path is similar to that proposed by Friedli and Velenik in the case of classical lattice systems (see Lemma 6.74 in Ref. Friedli and Velenik, 2017) but here applied to the mixed QM/MM case. In order to be able to use Jensen’s inequality, we will first need to show that the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ is indeed a probability measure. While for $\boldsymbol{p}$ it is obvious, it requires a little more attention for $\hat{{\boldsymbol{\rho}}}$ as it consists formally in a operator. Then, $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ will be written as an average value on probability measures, and Jensen’s inequality and strict convexity will do the rest.
.1 Rewriting $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ as an average value on probability measures
We use the adiabatic basis ${{\mathcal{B}}({\bf{Q}})=\{{|\Psi_{i}({\bf{Q}})\rangle}},0\leq i\leq\infty\}$ to develop the quantum or mixed quantum-classical operators. More precisely, ${\mathcal{B}}({\bf{Q}})$ consists in the complete set of eigenvectors of the following Schrodinger equation
$\big{(}\hat{{\bf h}}^{\text{tot}}_{\text{qm}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})\big{)}{|\Psi_{i}({\bf{Q}})\rangle}=E_{i}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle},$ (A.1)
such that the Hamiltonian $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is naturally written as
$\hat{{\bf h}}^{\text{tot}}_{\text{qm}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})=\sum_{i}E_{i}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|},$ (A.2)
and the operator $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ is written as follows
$\hat{{\boldsymbol{\rho}}}({\bf{Q}})=\sum_{i}{\boldsymbol{\rho}}_{ii}({\bf{Q}}){|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|}.$ (A.3)
Because of the properties of density matrices, the real-numbers ${\boldsymbol{\rho}}_{ii}({\bf{Q}})$ fulfill
${\boldsymbol{\rho}}_{ii}({\bf{Q}})\geq 0,\sum_{i\in{\mathcal{B}}({\bf{Q}})}{\boldsymbol{\rho}}_{ii}({\bf{Q}})=1.$ (A.4)
Therefore, as $\boldsymbol{p}({\bf{Q}})$ is a probability distribution and because of the properties of Eq. (A.4), if we define the function $\mu({\bf{Q}},i)$ as
${\mathbb{R}}^{3N_{{\text{qm}}}}\times{\mathbb{N}}\rightarrow{\mathbb{R}},$ (A.5)
$({\bf{Q}},i)\mapsto\mu({\bf{Q}},i)=\boldsymbol{p}({\bf{Q}}){\boldsymbol{\rho}}_{ii}({\bf{Q}}),$
then $\mu({\bf{Q}},i)$ fulfils the requirements for a probability measure, i.e.
$\mu({\bf{Q}},i)\geq 0\,\,\forall\,\,({\bf{Q}},i),\quad\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)=1.$ (A.6)
For the sake of compactness, we define ${\mathbf{X}}=({\bf{Q}},i)$, such that we can rewrite the probability measure as
$\mu({\bf{Q}},i)\equiv\mu({\mathbf{X}}),\quad\int\text{d}{\bf{Q}}\sum_{i}\equiv\int\text{d}{\mathbf{X}}.$ (A.7)
The equilibrium density $\hat{\mathcal{P}}_{\text{eq}}$ of Eq. (13) can be written using the adiabatic basis of Eq. (A.2) as follows
$\hat{\mathcal{P}_{\text{eq}}}({\bf{Q}})$ $=\frac{1}{Z}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}e^{-\beta\hat{{\bf V}}^{\text{tot}}_{\text{mm}}({\bf{Q}})}$ (A.8)
$=\frac{1}{Z}\sum_{i}e^{-\beta\epsilon({\bf{Q}},i)}{|\Psi_{i}({\bf{Q}})\rangle}{\langle\Psi_{i}({\bf{Q}})|},$
which corresponds to the measure $\mu({\mathbf{X}})=\mu_{\text{eq}}({\mathbf{X}})$ defined as follows
$\mu_{\text{eq}}({\mathbf{X}})\equiv\frac{e^{-\beta\epsilon({\mathbf{X}})}}{Z},$ (A.9)
where
$\epsilon({\mathbf{X}})=\epsilon({\bf{Q}},i)\equiv E_{i}({\bf{Q}})+\hat{{\bf V}}^{\text{tot}}_{\text{mm}}({\bf{Q}}).$ (A.10)
We introduce the notation for an average value over a probability measure $\mu({\mathbf{X}})$ as
$\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})f({\mathbf{X}})=\langle f\rangle_{\mu},$ (A.11)
such that the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ can be rewritten as an average value on the measure $\mu({\mathbf{X}})$
$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=$ $\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\big{(}\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}}\boldsymbol{p})\big{)}\}$ (A.12)
$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big{(}\epsilon({\bf{Q}},i)+\text{k}_{\text{B}}T\log(\mu({\bf{Q}},i))$
$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big{(}\epsilon({\mathbf{X}})+\text{k}_{\text{B}}T\log(\mu({\mathbf{X}}))$
$=$ $-\text{k}_{\text{B}}T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
$=$ $-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
$\equiv$ $F[\mu].$
According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ at $\mathcal{P}_{\text{eq}}$ leads straightforwardly to
$F[\mu_{\text{eq}}]=$ $-\text{k}_{\text{B}}T\log(Z),$ (A.13)
which is of course the correct Helmholtz free energy. Then two questions remain: i) is $-\text{k}_{\text{B}}T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
.2 Use of Jensen inequality to prove variational character
As the function $-\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer. The Jensen’s inequality can be stated as follows.
If $\varphi$ is a convex function (i.e. always above its tangent lines), $\mu({\mathbf{X}})$ a probability measure on ${\mathbf{X}}\in E$ where $E$ is a normed vector space, $g({\mathbf{X}})$ a function from $E$ to ${\mathbb{R}}$, then Jensen’s equality states that
$\varphi(\langle g\rangle_{\mu})\leq\langle\varphi(g)\rangle_{\mu},$ (A.14)
or written in terms of integrals
$\varphi\bigg{(}\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})g({\mathbf{X}})\bigg{)}\leq\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\varphi\big{(}g({\mathbf{X}})\big{)}.$ (A.15)
As $-\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
$g({\mathbf{X}})=\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})},$ (A.16)
$\varphi(x)=-\text{k}_{\text{B}}T\log(x).$
Therefore the inequality of Eq. (A.14) becomes
$-\text{k}_{\text{B}}T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
and as
$\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}$ $=\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\frac{e^{-\beta\epsilon(x))}}{\mu({\mathbf{X}})}$ (A.18)
$=Z,$
one can rewrite Eq. (A.17) as
$-\text{k}_{\text{B}}T\log(Z)\leq F[\mu],$ (A.19)
which guarantees the variational character. Eventually, the strict convexity of the $-\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into
$g({\mathbf{X}})=cst\Leftrightarrow\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})}=cst\Leftrightarrow\mu({\mathbf{X}})\propto e^{-\beta\epsilon({\mathbf{X}})},$ (A.20)
therefore implying that $\mu_{\text{eq}}({\mathbf{X}})$ is the only minimizer. We point out that this proof essentially consists in the Gibb’s inequality re-framed with convex analysis vocabulary.

### v5

@@
 Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures.
-However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations.
-This paper aims to develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation and serves as a starting point for further development.
+However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region
+This paper addresses the second point and develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation.
 We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble.
-We start with the adiabatic equilibrium density matrix for a system comprising $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles and, based on this, we propose a variational formulation of the Helmholtz free energy in terms of the full QM/MM density matrix.
-Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities, thus generalizing both cDFT and electronic DFT (eDFT) for QM/MM systems.
-We then reformulate the functional in order to explicitly make appear the eDFT and cDFT Levy-Lieb functionals, together with a new universal correlation functional for QM/MM systems.
+We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles.
+Then, we propose a variational formulation of the Helmholtz free energy in terms of the full , non-equilibrium, QM/MM density matrix.
+Taking advantage of permutational symmetry and thanks to constrained-search methods, we reformulate the computation of the Helmholtz free energy using only the quantum and classical one-body densities.
+Therefore, this paper generalizes both cDFT and electronic DFT (eDFT) to QM/MM systems.
+We then reformulate the functional to make the standard eDFT and cDFT Levy–Lieb functionals explicitly appear, together with a new universal correlation functional for QM/MM systems.
 A mean-field approximation is finally introduced in the context of solvation problems and we discuss its connection with several existing mixed cDFT-eDFT schemes.
+An extension to the semi-grand canonical ensemble, where the number of classical particles is allowed to fluctuate, is provided in the supplementary materials.
 I Introduction
@@
 Although electronic structure calculations at zero temperature can routinely handle systems with hundreds to even thousands of atoms, accounting for finite-temperature effects becomes crucial in many applications.
-Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion.
-In this context, it is natural to compute the forces acting on the nuclei at the quantum level and then to solve the classical equations of motion to propagate the trajectory.
-This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999).
-Despite its considerable achievements, it is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space.
+Such calculations must therefore be set in a statistical ensemble, where the leading-order entropic contribution arises from nuclear motion due to the typically large gaps in electronic energies.
+In this context, because of the large mass of nuclei with respect to that of electrons, it is physically relevant to maintain electrons at the quantum level and to treat nuclei as classical particles.
+Nevertheless, the forces acting on the nuclei are computed from the solutions of the quantum electronic ground state, i.e.
+from the solutions of the Born Oppenheimer Hamiltonian.
+This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
+Nevertheless, it is important to keep in mind that nuclear quantum effects play a significant role for light nuclei and can substantially alter the rates of certain chemical processes, such as proton transfer reactionsBała et al.
+(1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin, Robert, and Savéant (2010); Hammes-Schiffer and Stuchebrukhov (2010).
+Moreover, such proton transfers can also be strongly influenced by the solvation effects, specially when polar solvents are usedKapral (2015).
+In this context, two fundamentally different yet interconnected nuclear quantum effects are at work: (i) the intrinsic delocalization of nuclear wave packets due to the quantum nature of nuclei, and (ii) the nonadiabatic couplings that arise between various adiabatic electronic states.
+Nuclear delocalization can be captured through explicit nuclear wave functions (see, for example, Ref.
+Beck et al., 2000 for a review of time-dependent mean-field approaches, and Ref.
+Hammes-Schiffer, 2021 for the nuclear-electron orbital approach) or via path-integral formalisms such as ring-polymer molecular dynamicsHabershon et al.
+(2013).
+Nonadiabatic couplings, on the other hand, are commonly addressed using surface hopping methods based on classical nuclear trajectoriesTully (1990); Hammes-Schiffer and Tully (1994); Ben-Nun and Martı’nez (2000); Li et al.
+(2005); Richardson and Thoss (2013); Ananth (2013), modified time-dependent mean-field techniquesWorth et al.
+(2008), semiclassical expansions of the density matrix and its Wigner representationWorth et al.
+(2008), or even hydrodynamic formulations of quantum mechanics à la Bohm-de BroglieBurghardt and Cederbaum (2001a, b); Burghardt and Bagchi (2006).
+In contrast, the present work adopts a fully classical treatment of the solvent molecules and focuses on clarifying the formally exact framework for a given QM/MM partitioning, as typically addressed in ab initio molecular dynamics.
+Despite its considerable achievements, ab-initio molecular dynamics is computationally demanding due to the extensive number of QM calculations required to sample the exponentially large phase space.
 The convergence rate of the statistics can also vary significantly depending on the targeted properties and conditionsPan et al.
@@
 (2018).
-To alleviate ab-initio molecular dynamics computational burden, QM/MM methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM).
+To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland, Lyne, and Karplus (2000); Senn and Thiel (2009a); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM).
 The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion.
-However, even in QM/MM approaches, sampling the phase space of large systems remains computationally intensive.
+The choice of the classical model to represent molecular entities together with the choice of the QM/MM interaction lacks of a unique definition and consists into a whole field of researchRiniker (2018); Harrison et al.
+(2018); Unke et al.
+(2021); Wu et al.
+(2023); Böselt, Thürlemann, and Riniker (2021); Ho et al.
+(2025) involving both physically-motivated and machine-learning oriented approaches.
+However, despite the significant gain in computational cost with respect to ab initio molecular dynamics, sampling the phase space of large systems remains computationally intensive even in QM/MM approaches.
 Inspired by the extension of eDFT to QM systems at finite temperaturesMermin (1965), density-based formalisms have emerged for describing purely classical objects in statistical mechanics, particularly in the form of classical density functional theory (cDFT)Evans (1979).
@@
 Therefore, most equilibrium properties can be obtained through functional minimization over the classical one-body density rather than through tedious sampling of configurations in an exponentially large phase space.
-The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object—such as the wavefunction in QM or the probability distribution in MM—with the optimization of a functional that depends only on a low-dimensional quantity, i.e.
+The efficiency of both eDFT and cDFT stems from a common principle: each method replaces the challenge of computing a high-dimensional object, such as the wavefunction in QM or the probability distribution in MM, with the optimization of a functional that depends only on a low-dimensional quantity, i.e.
 the one-body density of either quantum electrons or classical molecules.
-The mathematical framework allowing for such a simplification is the so-called constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFT, although not usable in practical calculations.
+The mathematical framework allowing for such a simplification is the constrained-search principle introduced by Levy and LiebLevy (1979); Lieb (1983) which enables, under certain constraints of representativity, to give an explicit expression of the universal functionals of eDFT and cDFTDwandaru and Schmidt (2011), although not usable in practical calculations.
 Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al.
@@
 (2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
-In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians.However, the QM/MM interaction lacks of a unique practical definition.
-The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023).
+In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians.
+However, just as in molecular dynamics based QM/MM approaches, the coupling interaction between the QM and MM regions lacks of a unique practical definition.
+The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009b); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023).
 The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals.
@@
 Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al.
-(2007) justified their so-called joint density-functional formulation by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei.
-To derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei.
-Howvever, this step makes it difficult to pursue a rigorous derivation and their working functional is rather defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals.
+(2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem.
+The authors obtain a multi component electron/nuclei/solvent DFT formulation, which is justified by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei.
+However, to derive a QM/MM formulation, they treat the environment particles as classical by ’integrating out’ the electronic density associated with these nuclei, without explicitly indicating how this is done.
+Therefore, this step makes it difficult to pursue a rigorous derivation.
+Their working functional is instead defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals.
 In other workJeanmairet, Levesque, and Borgis (2020); Tang et al.
 (2020), a similar form of the functional was invoked without further theoretical justification.
-Thus, while QM/cDFT approaches offer potentially attractive computational scaling, their derivation from an exact theory remains somewhat ambiguous.
+Thus, while QM/cDFT approaches offer potentially attractive computational scaling and appears as a quite natural outcome, their derivation from an exact theory remains somewhat ambiguous.
 This work addresses this gap by establishing a rigorous mathematical framework that provides a DFT-equivalent description of mixed QM/MM systems.
@@
 Extending this formulation to the grand canonical ensemble for classical particles is conceptually straightforward but introduces additional complexity in the notations due to the varying number of particles.
-The main difference between the two ensembles is that, in the grand canonical ensemble, there exists a unique correspondence between the equilibrium density and the external potential.
-In contrast, in the canonical ensemble, the equilibrium density corresponds to a family of external potentials, defined up to an additive constantLutsko (2022).
-Within the present paper, the foundational object from which the QM/MM DFT formulation is derived is the explicit expression of the adiabatic equilibrium density of a mixed QM/MM system in the canonical ensemble.
-This expression, which was previously established in the adiabatic basis by Nielsen et.
+We therefore do not include it in the core of the paper, but rather in the supplementary materials.
+This derivation follows a similar route to its formally lighter canonical counterpart, which is reported in the present paper.
+The main difference between the two ensembles is the existence of a unique mapping between the equilibrium density and the external potential in the grand canonical ensemble.
+In contrast, in the canonical ensemble, the equilibrium density is associated to a family of external potentials, defined up to an additive constantGonzález et al.
+(1997); Lutsko (2022).
+Our derivation begins with two key assumptions: (a) the entire system is partitioned into a QM region and a MM region; and (b) specific models are available for both the MM subsystem and the QM/MM coupling.
+As such, we do not address the QM/MM crossover or the accuracy of the MM and QM/MM models themselves.
+We assume these models are given, as is standard in practical QM/MM calculations.
+Under these assumptions, the primary objective of this paper is to derive, in a straightforward manner, the one-body density functional theory (DFT) formulation for a QM/MM system.
+Therefore, The key steps of this derivation are as follows: i) We begin by defining the general Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian (see Sec.
+II.1).
+ii) We then express the exact solution of the problem using full-system quantities (i.e., those describing all particles).
+This includes the Helmholtz free energy of the QM/MM system, along with the corresponding equilibrium density matrix and partition function (see Sec.
+II.2).
+iii) Next, we reformulate the problem as the minimization of a functional over generic full-system density matrices.
+In our case, this means expressing the Helmholtz free energy as a functional of QM/MM density matrices that describe all particles in the system (see Sec.
+II.3.1).
+iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.3.2).
+v) Under certain assumptions about the form of the external potentials, we show how the external part of the functional can be written as a linear functional of one-body densities.
+In our case, this results in a functional of both classical and quantum one-body densities (see Sec.
+III.1).
+vi) Leveraging the many-to-one correspondence between full-system and one-body quantities, we perform a Levy–Lieb constrained search to rewrite the intrinsic free energy functional as a universal functional of the one-body densities (see Sec.
+III.1).
+vii) We then establish a connection between electronic DFT (eDFT) and classical DFT (cDFT) by introducing a universal QM/MM correlation functional of the one-body QM and MM densities (see Sec.
+III.2).
+viii) Finally, we discuss this formulation in the context of solvation problems and recent developments in QM/cDFT coupling approaches (see Sec.
+III.3).
+The Appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ contains a proof of the variational properties of the QM/MM functional introduced here, and the Supplementary Material provides a recall of the Von-Neuman variational principle in statistical QM for the canonical ensemble, an alternative derivation of the QM/MM equilibrium density previously established in the adiabatic basis by Nielsen et.
 al.
-in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), is recovered here following an alternative route (see Sec.
-II of the supplementary materials).
-Building on this, we propose a functional formulation of the free energy in terms of general QM and MM high-dimensional densities (see Sec.
-Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ for a proof of the variational properties).
-By applying the Levy-Lieb constrained search and assuming additive two-body QM/MM interactions, we then naturally express the free energy in terms of the one-electron density and the classical one-body density.
-We explicitly establish the connection between eDFT and cDFT through the introduction of a new universal QM/MM correlation one-body density functional.
-Additionally, we discuss a mean-field approximation and its relationship with previously proposed cDFT/eDFT couplings in the context of QM/MM solvation problems.
-We also note that this work can be seen as the QM/MM and finite-temperature counterpart of the recently published study by Fromager and LasorneFromager and Lasorne (2024), which formulated an exact DFT framework incorporating both electrons and nuclei at the quantum level—going beyond the Born-Oppenheimer approximation.
-The paper is organized as follows.
-In Sec.
-II, we begin by discussing the adiabatic equilibrium density for a mixed QM/MM system in the canonical ensemble (Sec.
-II.1 and II.2).
-Then, in Sec.
-II.3, using the explicit form of the equilibrium density, we reformulate the computation of the free energy as a functional minimization over general high-dimensional mixed QM/MM densities.
-Sec.
-III focuses on expressing this problem explicitly in terms of the one-body QM and MM densities.
-In Sec.
-III.1, we employ the Levy-Lieb constrained search to reformulate the high-dimensional QM/MM functional problem by recognizing a many-to-one correspondence between the high-dimensional QM/MM quantities and the one-body QM and MM densities.
-In Sec.
-III.2, we establish a connection between eDFT and cDFT by introducing a universal QM/MM correlation functional of the one-body QM and MM densities.
-Sec.
-III.3 then discusses this formulation in the context of solvation problems and recent QM/cDFT coupling approaches.
-Finally, Sec.
-IV summarizes the main results and conclusions.
-The supplementary materials provide an alternative derivation of the equilibrium density for a QM/MM system that does not rely on Liouvillian operators, and the appendix Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ provides a proof of the variational properties of the QM/MM functional.
+in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), and an extension to the semi grand canonical ensemble of the present formalism.
+As this work is intended for both the QM and MM communities, which often differ in terminology and conceptual frameworks, we highlight certain elements throughout the text to aid accessibility.
+Some of these clarifications may appear self-evident to some readers; we hope, however, that they do not unduly burden the readability of the article.
 II Quantum-classical system in the Canonical ensemble: equilibrium density and variational formulation
-The main goal of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables.
+The objective of this section is to give a variational formulation for the Helmholtz free energy of a mixed QM/MM system based on a functional minimization over general densities involving $N_{{\text{qm}}}$ and $N_{{\text{mm}}}$ variables.
+This is essentially a QM/MM equivalent to the Gibbs variational principle in classical statistical physics or the Von-Neumann variational principle in quantum statistical physics (see the supplementary information for a short review).
+This variational principle, although hardly usable in practice due to the high-dimensionality of the objects it manipulates, is an important milestone to properly obtain the one-body formulation, which is the aim of this paper.
 We define in Sec.
-II.1 our working Hamiltonian describing the QM/MM system.
+II.1 the working Hamiltonian that describes the QM/MM system.
+Again, we do not discuss the validity of the partitioning or the quality of the coupling term, we assume that these preliminary considerations have already been addressed.
 The equilibrium density of a mixed QM/MM system is then given in Sec.
@@
 II.1 The Hamiltonian of the quantum-classical system
-We adopt the following notation conventions : i) lowercase letters will refer to purely quantum objects (both variables and operators) while capital letters will refer to either purely classical or mixed quantum-classical objects and ii) all bold letters refer to collective variables or operators (i.e.
-relative to all quantum or classical particles) and non bold letters will refer to individual or pairwise variables or operators.
-The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$, and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
-The Hamiltonian of the total system is written as
+We consider a QM/MM system, where the partition and the choice of the model to describe the MM part and the coupling term have already been made
+We adopt the following notation conventions: (i) lowercase letters denote purely quantum quantities (variables or operators), while uppercase letters denote classical or mixed quantum–classical quantities; (ii) bold symbols denote collective quantities (over all particles), while non-bold symbols denote single-particle or pairwise quantities.
+The only exception is the number of quantum particles, labeled $N_{{\text{qm}}}$.
+Atomic units are used throughout.
+The total system is divided into a set of $N_{{\text{mm}}}$ identical classical particles, described by their classical positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and linear momentum ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$ and a set of $N_{{\text{qm}}}$ identical quantum particles described by their position variables ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$.
+The total Hamiltonian is written as the sum of the QM and MM Hamiltonians, and a QM/MM coupling term, namely
 $\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P},{\bf{q}})=\hat{\bf h}_{{\text{qm}}}({\bf{q}})+{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}}).$ (1)
 In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous.
-In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons) in interaction with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$ local in position representation, and interacting among themselves with a pairwise potential ${w}_{{\text{qm}}}({{q}}_{1},{{q}}_{2})$.
+In Eq.(1), $\hat{\bf h}_{{\text{qm}}}$ is the quantum Hamiltonian operator describing $N_{{\text{qm}}}$ quantum identical particles of mass $m$ (typically electrons).
+These particles interact with an external potential $v^{\text{ext}}_{{\text{qm}}}({{q}})$, which is local in the position representation, and with each other through a pairwise potential ${w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j})$.
 The explicit form of $\hat{\bf h}_{{\text{qm}}}$ reads then
@@
 with
-$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\hbar\nabla_{{{q}}_{i}},$ (3)
-$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{j}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
-Then, ${\bf H}_{{\text{mm}}}$ is the classical Hamiltonian function representing $N_{{\text{mm}}}$ identical classical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves with a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$.
-These particles are typically effective particles such as atoms or rigid molecules instead of the individual electrons composing them and they interact through model, parametrized, force fields.
+$\hat{\bf t}=\sum_{i=1}^{N_{{\text{qm}}}}\frac{\hat{p}_{i}^{2}}{2m},\qquad\hat{p}_{i}=-i\nabla_{{{q}}_{i}},$ (3)
+$\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}=\sum_{i=1}^{N_{{\text{qm}}}}v^{\text{ext}}_{{\text{qm}}}({{q}}_{i}),\qquad\hat{\bf w}_{{\text{qm}}}=\sum_{j=1}^{N_{{\text{qm}}}}\sum_{i>j}{w}_{{\text{qm}}}({{q}}_{i},{{q}}_{j}).$ (4)
+Then, ${\bf H}_{{\text{mm}}}$ is the classical model Hamiltonian function representing $N_{{\text{mm}}}$ identical particles of mass $M$, in interaction with a local external potential $V^{\text{ext}}_{{\text{mm}}}({{Q}})$ and interacting among themselves through a local pairwise potential ${W}_{{\text{mm}}}({{Q}}_{i},{{Q}}_{j})$.
+These are typically effective particles, such as atoms or rigid molecules, and they interact with each other through parametrized force fields.
 The explicit form of the classical Hamiltonian ${\bf H}_{{\text{mm}}}$ reads then
@@
 $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}=\sum_{i=1}^{N_{{\text{qm}}}}\sum_{j=1}^{N_{{\text{mm}}}}{W}_{{\text{qm}}}^{{\text{mm}}}({{q}}_{i},{{Q}}_{j})$ (8)
-The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position ${Q}$ and a fundamental particle (such as an electron) at position ${q}$.
-II.2 Mixed quantum-classical equilibrium density
-The equilibrium density corresponding to the Hamiltonian of Eq. (1) can be written as
-$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}},$ (9)
-where $\beta=(\text{k}_{\text{B}}T)^{-1}$, $\text{k}_{\text{B}}$ is the Boltzmann constant, and the dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables ${\bf{Q}}$ is here explicitly highlighted.
-This result is nothing but the real-space representation of that obtained by Nielsen et.
-al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis.
-The result of Eq.
-(9) can be obtained by starting from an Hamiltonian where both sets of particles are considered to be quantum.
-Then, one performs a Wigner transform over the $N_{{\text{mm}}}$ particles only and eventually a semi-classical limit is obtained by letting the ratio of masses $(m/M)^{1/2}$ tends towards 0.
-Conserving only the zeroth order term of this semi-classical limit yields to Eq.
-(9).
-This expansion in terms of the ratio of the masses is natural since the quantum particles, typically electrons, are usually much lighter than the classical ones, typically atoms or molecules.
-The original derivation of Nielsen et.
-al.Nielsen, Kapral, and Ciccotti (2001) used the Liouvillian imaginary time evolution equation of the density matrix together with the Poisson-brakets operatorsİmre et al.
-(1967), which is a well-known formalism in the quantum dynamics communityKapral and Ciccotti (1999); Nielsen, Kapral, and Ciccotti (2001); Kapral (2006, 2015).
-However, we provide in the supplementary materials an alternative derivation which involves a similarity transformation and resorts to the Baker-Campbell-Hausdorf expansion to express the Wigner transform Hamiltonian as a series of nested commutators, which might appear more familiar to the ”static” quantum chemistry community.
-In Eq. (9), the normalization is the partition function of the QM/MM system which is expressed as
-$Z^{\text{tot}}={\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}}\int\text{d}{\bf{Q}}\text{d}{\bf P}\!\!\int\!\text{d}{\bf{q}}$ ${\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}$ (10)
-$e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}.$
-The corresponding Helmholtz free energy is given by
-$F_{0}^{\text{tot}}=-\text{k}_{\text{B}}T\log(Z^{\text{tot}}).$ (11)
-While the expressions of Eqs.
-(9) and (10) are given in real-space, they can easily be written in the adiabatic basis, where one explicitly sees that the density matrix is diagonal as found by Nielsen et.
+The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position Q{Q} and a fundamental particle (such as an electron) at position q{q}.
+II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
+We start by defining the exact equilibrium density matrix (EDM) of the QM/MM system in the canonical ensemble.
+This quantity is the analogue of the equilibrium density matrix of a purely quantum system or, equivalently, of the equilibrium probability density of a purely classical (MM) system.
+With the Hamiltonian of Eq. (1), the EDM takes the following form
+$\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ $=\frac{{\langle{\bf{q}}|}e^{-\beta\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P})}{|{\bf{q}}^{\prime}\rangle}}{Z^{\text{tot}}}$ (9)
+$=\frac{{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}^{\prime}\rangle}e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}}{Z^{\text{tot}}}.$
+In Eq.
+(9) $\beta=T^{-1}$ in atomic units, and $|{\bf{q}}\rangle$} and $|{\bf{q}}^{\prime}\rangle$} are position eigenstates for the $N_{{\text{qm}}}$ quantum variables.
+The dependency of $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}$ on the classical variables, $\bf{Q}$}, has been made explicit.
+Because Eq.
+(9) involves the exponential of a quantum operator, the QM/MM EDM is itself a quantum operator, namely a quantum density matrix (i.e.
+positive definite and of trace unity).
+However, it is parametrized by the classical variables $({\bf{Q}},{\bf P})$ and therefore the QM system depends on the configuration of the classical particles.
+To understand the physical meaning of the EDM, it is instructive to analyse its diagonal elements, i.e.
+when ${\bf{q}}={\bf{q}}^{\prime}$.
+The quantity $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}})$ gives the probability of finding the quantum particles at positions ${\bf{q}}=\{{{q}}_{1},\ldots,{{q}}_{N_{{\text{qm}}}}\}$ while the classical particles have positions ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$ and momenta ${\bf P}=\{{{P}}_{1},\ldots,{{P}}_{N_{{\text{mm}}}}\}$.
+Off-diagonal elements encode purely quantum effects, such as nonlocality, and are necessary to compute the action of quantum operators—for example, the kinetic energy operator, which involves a Laplacian in the quantum subsystem.
+A pictorial illustration of the physical meaning of the equilibrium density of Eq.
+(9) is provided by the Born-Oppenheimer description of a molecule, i.e.
+QM electrons and classical nuclei.
+In that case, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, with no external potential for either the QM or MM parts.
+The operator $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ represents the nuclei-electron attraction, while ${\bf W}_{{\text{mm}}}({\bf{Q}})$ represents the Coulomb repulsion between the nuclei.
+Thus, $\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P})$ entering in the exponential of Eq.
+(9) corresponds the Born-Oppenheimer Hamiltonian.
+From a more formal point of view, it is worth noticing that Eq.
+(9) corresponds to the real-space representation of the EDM, originally obtained by Nielsen et.
+al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis, i.e.
+in the basis of eigenvectors of $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$.
+In this basis, the density matrix is diagonal , as found by Nielsen et.
 al.Nielsen, Kapral, and Ciccotti (2001).
+In the case of quantum electrons and classical nuclei, the adiabatic states are the eigenstates of the Born-Oppenheimer Hamiltonian.
+Therefore, Eq.
+(9) ignores the non-adiabatic couplings between these states.
+In the context of ab initio MD, this implies that $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ can be obtained from classical nuclear trajectories, with forces averaged over all adiabatic electronic states weighted by the corresponding Boltzmann factors.
+Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})$ is referred to as the adiabatic QM/MM density matrix.
+There is a flourishing literature addressing methods going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM systems (see for instance Ref.
+Kapral, 2015 for a review), but treatment of non adiabatic couplings is beyond the scope of this paper.
+[…]
+In Eq. (9), the normalization corresponds to the partition function of the QM/MM system which is expressed as
+$Z^{\text{tot}}=\frac{1}{N_{{\text{mm}}}!}$ $\int\text{d}{\bf{Q}}\text{d}{\bf P}\,e^{-\beta{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})}$ (10)
+$\int\!\text{d}{\bf{q}}{\langle{\bf{q}}|}{e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}}{|{\bf{q}}\rangle}.$
+The corresponding Helmholtz free energy is therefore given by
+$F_{0}^{\text{tot}}=-T\log(Z^{\text{tot}}).$ (11)
+Any equilibrium property $\langle O\rangle_{\text{eq}}$ can be obtained from derivatives of $Z^{\text{tot}}$.
+Alternatively, $\langle O\rangle_{\text{eq}}$ can be computed as the trace of the product of $\mathcal{M}$ and the operator $\hat{O}$, which may depend on both quantum and classical variables
+$\langle O\rangle_{\text{eq}}$ $=\int\text{d}{\bf{Q}}\text{d}{\bf P}\,\int\!\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}\hat{O}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}}^{\prime},{\bf{q}})$ (12)
+$\equiv\text{Tr}\{\hat{O}\hat{\mathcal{M}}\}$.
 Although the expression of the equilibrium density in Eq.
-(9) is relatively simple, it calls some comments.
+(9) is relatively simple, it benefits from further comments.
 Firstly, as usual in classical statistical mechanics, the contribution to the Helmholtz free energy coming from the purely classical kinetic term in ${\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})$ can be computed analytically.
-Therefore it is worth splitting the equilibrium density in terms of ${\bf{Q}}$ and ${\bf P}$ dependence
+Therefore, it is convenient to separate the equilibrium density according to its dependence on $\bf{Q}$} and $\bf P$}
 $\mathcal{M}({\bf{Q}},{\bf P},{\bf{q}},{\bf{q}}^{\prime})=\frac{e^{-\beta{\bf T}}}{z_{{\bf P}}}\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}),$ (13)
@@
 ${{\bf{V}}}_{\text{mm}}$ $={\bf V}^{\text{ext}}_{{\text{mm}}}+{\bf W}_{{\text{mm}}},$ (15)
-$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!h^{3N_{{\text{mm}}}}}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg{(}\frac{2\pi M}{\beta h^{2}}\bigg{)}^{\frac{3}{2}N_{{\text{mm}}}},$ (15)
+$z_{{\bf P}}$ $=\frac{1}{N_{{\text{mm}}}!}\int\text{d}{\bf P}e^{-\beta{\bf T}}=\frac{1}{N_{{\text{mm}}}!}\bigg(\frac{2\pi M}{\beta}\bigg)^{\frac{3}{2}N_{{\text{mm}}}},$ (16)
 $Z$ $=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}{\langle{\bf{q}}|}e^{-\beta(\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}}))}{|{\bf{q}}\rangle}e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})},$ (17)
@@
 where
-$F_{{\bf P}}=-\text{k}_{\text{B}}T\log(z_{{\bf P}}),$ (19)
+$F_{{\bf P}}=-T\log(z_{{\bf P}}),$ (20)
 and
-$F_{0}=-\text{k}_{\text{B}}T\log(Z)$ (20)
-is the interacting part of the Helmholtz free energy of the QM/MM system.
-Secondly, the object $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ in Eq.(13) can be expressed as the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
+$F_{0}=-T\log(Z).$ (21)
+$F_{0}$ represents the contribution to the Helmholtz free energy of the QM/MM system arising from particle interactions.
+Computing $F_{0}$ from Eq.
+(21) is infeasible for realistic systems, as it requires evaluating Eq.
+(17).
+The aim of this paper is to develop an alternative, more practical, route to $F_{0}$.
+The EDM $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ of Eq.(14) can be decomposed into the product of a purely quantum density matrix ${\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})$ (more precisely its matrix element evaluated in real space) and a purely classical function $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$, i.e.
 $\mathcal{P}_{\text{eq}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})={\boldsymbol{\rho}}_{\text{eq}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}})\boldsymbol{p}_{\text{eq}}({\bf{Q}}),$ (22)
@@
 $Z$ $=\int\text{d}{\bf{Q}}Z^{{\text{qm}}}({\bf{Q}})e^{-\beta{{\bf{V}}}_{\text{mm}}({\bf{Q}})}.$ (27)
-It is noteworthy that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ defined in Eq.
-(22),(23) is a positive definite operator and that $\boldsymbol{p}_{\text{eq}}({\bf{Q}})$ of Eq.
-(25) is also positive definite and normalized to unity.
-Therefore, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are simply a special case of more general density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively.
-This will allow us in Sec.
-II.3 to propose a variational formulation of the Helmholtz free energy for a QM/MM system.
-Additionally, this decomposition of the equilibrium density as a product of a purely quantum equilibrium density and a purely classical equilibrium density could lead to the somehow strange conclusion that there is no correlation between the quantum and classical systems, even if the two systems interact.
-Nevertheless, one must pay attention to the fact that the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$.
-Therefore, the two systems are indeed correlated.
-Thirdly, an obvious illustration of the meaning of the equilibrium density of Eq.
-(13) is that corresponding to the Born-Oppenheimer approximation.
-In the latter case, the light particles described by ${\bf{q}}$ are electrons and the heavy particles described by ${\bf{Q}}$ are the classical nuclei.
-There, $\hat{\bf h}_{{\text{qm}}}$ is composed of the kinetic operator of the electrons and their mutual coulomb interaction, there is no external potential both for the QM and MM parts, $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$ is the nuclei-electron attraction, and ${\bf W}_{{\text{mm}}}({\bf{Q}})$ is the Coulomb repulsion between the nuclei, such that the operator happening in the exponential of Eq.
-(13) is the Born-Oppenheimer Hamiltonian.
-A follow up on this comment is that, as mentioned earlier, $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is diagonal in the adiabatic basis, which is equivalent to say that non-adiabatic couplings are completely neglected.
-In the context of ab initio MD, this implies that $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ can be obtained by averaging, over all possible adiabatic electronic states, the classical trajectories of nuclei computed from the forces of a given adiabatic electronic state.
-Consequently, it entirely ignores surface hopping between different adiabatic states, which is why $\hat{{\boldsymbol{\rho}}}_{\text{eq}}({\bf{Q}})$ is referred to as the adiabatic QM/MM density matrix.
-There is a flourishing literature concerned with going beyond this adiabatic approximation, specially in the context of quantum dynamics or statistical mechanics for QM/MM treatments (see for instance Ref.
-Kapral, 2015 for a review), but the treatment of non adiabatic couplings goes beyond our QM/MM scheme.
-II.3 Reformulation of the Helmholtz free energy for a QM/MM system within a variational framework
+It is noteworthy that, for each value of $\bf{Q}$}, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}({\bf{Q}})$ defined in Eq.
+(23) and (24) is a density matrix, i.e.
+positive definite and of unity norm.
+Similarly, $\boldsymbol{p}_{\text{eq}}$ of Eq.
+(26) is a probability distribution.
+Therefore, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$ and $\boldsymbol{p}_{\text{eq}}$ are special cases of more general objects: density matrices representing $N_{{\text{qm}}}$ quantum particles and probability distribution functions representing $N_{{\text{mm}}}$ classical particles, respectively.
+This will allow to formulate a variational formulation of the Helmholtz free energy for a QM/MM system in Sec.
+II.3.
+Last but not least, the decomposition of the equilibrium density of Eq.
+(22) into a product of a purely quantum equilibrium density and a purely classical equilibrium density might suggest that the quantum and classical subsystems are uncorrelated, even though they interact.
+Nevertheless, the quantum part of the equilibrium density, $\hat{{\boldsymbol{\rho}}\,}_{\text{eq}}$, depends on the coordinates of the classical systems through the QM/MM interaction potential $\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$.
+The classical part, $\boldsymbol{p}_{\text{eq}}$, is in turn affected by the quantum degrees of freedom through its normalization, $Z^{{\text{qm}}}({\bf{Q}})$.
+To see this more explicitly, one can notice that the form of Eq.
+(26) can be rewritten as
+${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\boldsymbol{p}_{\text{eq}}({\bf{Q}})=\frac{1}{Z}e^{-\beta\big({{\bf{V}}}_{\text{mm}}({\bf{Q}})+U_{\text{eff}}({\bf{Q}})\big)},}$ (28)
+where now the classical particles experience an additional the effective potential $U_{\text{eff}}({\bf{Q}})$ due to the presence of the QM system, which is defined as
+$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U_{\text{eff}}({\bf{Q}})=-\frac{1}{\beta}\log\big(Z^{{\text{qm}}}({\bf{Q}})\big).$} (29)
+Therefore, the two systems are correlated.
+II.3 Variational formulation of the Helmholtz free energy for a QM/MM system
+Both classical and quantum DFT can be formulated through the Levy-Lieb constrained-search formalism.
+The first step is to express the quantum ground state energy and the classical Helmholtz free energy as a general functional minimization problem, defined over the many-body functions, namely the electronic density matrix and the probability distribution, respectively.
+Since the goal of this work is to obtain a DFT formulation of the QM/MM problem, the first step is to reformulate the Helmholtz free energy of Eq.
+(21) as a functional minimization problem over QM/MM density matrices whose form generalize the EDM of Eq.
+(22).
+This is the aim of this section.
+To the best of our knowledge, this was never formalized before.
 II.3.1 Variational formulation
-As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we want to reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq.
+As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we can now reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq.
 (22) is the minimizer of that functional and that it yields the correct Helmholtz free energy of Eq.
-(20).
-Moreover, since the ultimate goal is to formulate a variational principle involving both quantum and classical densities, it is natural to use the decomposition of Eq.
 (21).
@@
 $F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (30)
-where the operator $\hat{{\boldsymbol{\rho}}}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
+where the operator $\hat{{\boldsymbol{\rho}}\,}$ is a quantum density matrix representing $N_{{\text{qm}}}$ particles but also parametrically depending on all classical variables ${\bf{Q}}=\{{{Q}}_{1},\ldots,{{Q}}_{N_{{\text{mm}}}}\}$, i.e.
 $\hat{{\boldsymbol{\rho}}\,}\equiv\hat{{\boldsymbol{\rho}}\,}({\bf{Q}}),$ (31)
 and $\boldsymbol{p}$ is a classical probability density representing $N_{{\text{mm}}}$ classical particles and not depending on any quantum variable.
-Therefore, ${\boldsymbol{\rho}}$ and $\boldsymbol{p}$ must be normalized to be density matrices and probability measures, respectively, i.e
+Therefore, $\hat{{\boldsymbol{\rho}}\,}$ and $\boldsymbol{p}$ must be normalized to unity in order to be acceptable density matrices and probability measures, respectively, i.e
 $\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})=1\quad\forall\,\,{\bf{Q}},$ $\int\text{d}{\bf{Q}}\,\,\boldsymbol{p}({\bf{Q}})=1,$ (32)
@@
 such that that the product of the two functions is normalized to unity
-$\text{Tr}\{\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\}=\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (31)
+$\int\text{d}{\bf{Q}}\int\text{d}{\bf{q}}\,\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})=1.$ (34)
 Therefore, the equilibrium density of the mixed QM/MM system described in Eq.
@@
 As shown in the appendix (see section Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$), if we define the Helmholtz functional as follows
-$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}],$ (32)
-where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has was accounted for in $F_{{\bf P}}$)
+$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\mathcal{U}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (35)
+where $\mathcal{U}$ is an energy functional excluding the classical kinetic term (which has already been accounted for in $F_{{\bf P}}$)
 $\mathcal{U}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\text{Tr}\big\{\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}(\hat{\bf h}_{{\text{qm}}}+\hat{{\bf v}}^{\text{ext}}_{{\text{qm}}}+{\bf W}_{{\text{mm}}}+{\bf V}^{\text{ext}}_{{\text{mm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}})\big\},$ (36)
@@
 then the only minimizer of the functional of Eq. (35) is $(\hat{{\boldsymbol{\rho}}\,}_{\text{eq}},\boldsymbol{p}_{\text{eq}})$ of Eq. (22), i.e
-$-\text{k}_{\text{B}}T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}},\boldsymbol{p}),$ (35)
+$-T\log(Z)\leq F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\quad\forall\,\,(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}),$ (38)
 and
-$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=F_{0}=-\text{k}_{\text{B}}T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (36)
+$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=F_{0}=-T\log(Z)\Leftrightarrow(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})=(\hat{{\boldsymbol{\rho}}\,}_{\text{eq}},\boldsymbol{p}_{\text{eq}}).$ (39)
+This variational formulation, is the analogue of the Gibbs or Von Neumann variational principles, but applied to a QM/MM system.
+As with any variational formulation, it naturally enables the construction of approximations by imposing constraints on the variables of minimization, the archetypal example being the mean-field approach, which is briefly discussed in Sec.
+III.3.
 II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
-Because $\boldsymbol{p}$ does not depend on the quantum variables and that we chose the normalization condition of Eq. (29), we can naturally decompose the entropy as
+Because the function $\boldsymbol{p}$ does not explicitly depend on the quantum variables and because we chose the normalization conditions of Eq. (32), we can naturally decompose the entropy as
 $\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}],$ (40)
-with $S_{{\text{mm}}}[\boldsymbol{p}]$ being the purely classical dimension-less entropy
+where $S_{{\text{mm}}}[\boldsymbol{p}]$ is the purely classical dimension-less entropy
 $S_{{\text{mm}}}[\boldsymbol{p}]$ $=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\log(\boldsymbol{p}({\bf{Q}})),$ (41)
 and $S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is the quantum-classical dimension-less entropy
-$S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=-\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\log(\hat{{\boldsymbol{\rho}}})\big{\}}$ (39)
+[…]
 $=-\int\text{d}{\bf{Q}}\,\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\,{\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})\log({\boldsymbol{\rho}}({\bf{q}},{\bf{q}},{\bf{Q}})),$
@@
 $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})\int\text{d}{\bf{q}}\int\text{d}{\bf{q}}^{\prime}{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime}){\boldsymbol{\rho}}({\bf{q}}^{\prime},{\bf{q}},{\bf{Q}}).$
-It is noteworthy here to highlight that the form of Eq.
-(45) is quite general as it consists in a priori in potential with a non local kernel ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$.
-This form of interaction includes for instance Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
+It is worth noting that the form of Eq.
+(48) is quite general, since ${\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}},{\bf{q}}^{\prime})$ is a priori non local.
+This form of interaction includes, for example, Lennard-Jones potentials, frozen density Coulomb interaction or density overlap models for Pauli repulsionWheatley and Price (1990).
 Eventually, once all terms are inserted in Eq. (30), the Helmholtz free energy $F_{0}$ can be obtained as
-[…]
-$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\big{(}S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big{)}\big{\}}.$
+$F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}\big\{$ $E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ (49)
+$+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\big(S_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+S_{{\text{mm}}}[\boldsymbol{p}]\big)\big\}.$
 Eq.
-(46) is therefore a variational formulation of $F_{0}$ for a mixed QM/MM system in terms of a minimization over the couple $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$, which, to the best of our knowledge, was never formalized before.
+(49) therefore provides a variational formulation of $F_{0}$ for a mixed QM/MM system, expressed as a minimization over the couple $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$.
 One can see from Eqs.
-(38), (41) and (42) that the functionals corresponding to the MM part of the energy and entropy are completely independent of the quantum part of the system, as they depend only on the classical probability $\boldsymbol{p}$.
-Nevertheless, because of the term $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$, the quantum and classical parts are coupled in the minimization process of Eq.
-(46).
-Also, one can see that the functionals corresponding to the QM parts of the energy and entropy explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
+(41), (44) and (45) that the functionals corresponding to the MM part of the energy and entropy (i.e.
+$E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$, $E_{{\text{mm}}}[\boldsymbol{p}]$ and $S_{{\text{mm}}}[\boldsymbol{p}]$) are, per se, completely independent of the quantum part of the system.
+They only depend on the classical probability $\boldsymbol{p}$.
+Nevertheless, the quantum and classical parts remain coupled in the minimization of Eq.
+(49) through the coupling functional $E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$.
+In contrast, the QM parts energy and entropy functionals explicitly depend on the classical MM probability distribution $\boldsymbol{p}$.
 Although Eq.
-(46) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the variables $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ involved in the minimization are nevertheless very high-dimensional.
-Therefore, the goal of the following sections is to reformulate the variational principle of Eq.
-(46) using functionals of significantly lower-dimensional densities, i.e.
+(49) provides a variational formulation of the Helmholtz free energy for a QM/MM system, the minimization variables $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ remain very high-dimensional because they explicitly describe every individual particle in the system.
+The objective of the following section is to reformulate the variational principle of Eq.
+(49) using functionals depending on much lower-dimensional objects, i.e.
 the one-body QM and MM densities.
 III One-body density formulation through Levy-Lieb constrained search formulation
-An important consequence of the one-body nature of both $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, along with the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}}$ is that the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ in Eqs.
-(42) and (44) can be computed using one-body objects.
-Therefore, by applying the Levy-Lieb constrained search, we can reformulate the variational problem of Eq.
-(46) in terms of a functional that depends only on the one-body quantum and classical densities, rather than the full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables.
+Starting from the variational formulation of the Helmholtz free energy for the QM/MM system in terms of the $N_{{\text{qm}}}+N_{{\text{mm}}}$ quantities in Eq.
+(49), we want to obtain a one-body DFT formulation.
+As in classical and quantum DFT, the key idea is to split the total energy into an external part and an intrinsic part.
+More precisely, because of the one-body nature of both classical and quantum external potentials $V^{\text{ext}}_{{\text{mm}}}({Q})$ and $v^{\text{ext}}_{{\text{qm}}}({q})$, and the permutational symmetry of both $\boldsymbol{p}$ and $\hat{{\boldsymbol{\rho}}\,}$, the corresponding energy terms $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ and $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ in Eqs.
+(45) and (47) can be expressed in terms of the one-body densities.
+Then, by applying the Levy-Lieb constrained search, we can also reformulate the intrinsic part of the energy as a functional of the one-body quantum and classical densities.
+Therefore, the variational problem of Eq.
+(49) is directly expressed as a functional of the one-body QM and MM densities rather than in terms of full $N_{{\text{mm}}}+N_{{\text{qm}}}$ variables quantities.
 III.1 Classical and quantum one-body density and the Levy-Lieb constrained search
-We naturally introduce the one-body classical density $n({Q})$ as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq.
+The classical one-body density is naturally defined as the trace over all quantum variables and all classical variables but one of the product $\hat{{\boldsymbol{\rho}}\,}({\bf{q}},{\bf{Q}})\boldsymbol{p}({\bf{Q}})$, which, thanks to the normalization of Eq.
 (32), is simply
@@
 […]
-It is noteworthy from Eq.
-(48) that, because of the ${\bf{Q}}$ dependence in $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$, the classical $N$-body probability $\boldsymbol{p}({\bf{Q}})$ contributes to the one-body quantum density $\rho({q})$, whereas from Eq.
-(47) it is clear that the quantum $N$-body density matrix $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ does not contribute to the classical one-body density.
-By using permutation symmetry of both $\hat{{\boldsymbol{\rho}}}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to exchange of either QM or MM variables, we can then write the two terms arising from the external potential as functionals of the QM and MM one-body densities
+Thanks to the permutational symmetry of both $\hat{{\boldsymbol{\rho}}\,}({\bf{Q}})$ and $\boldsymbol{p}({\bf{Q}})$ with respect to the exchange of either QM or MM coordinates, the two external functionals can be rewritten as functionals of the QM and MM one-body densities, respectively
 $E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]$ $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}}){\bf V}^{\text{ext}}_{{\text{mm}}}({\bf{Q}})$ (52)
@@
 $\equiv(v^{\text{ext}}_{{\text{qm}}}|\rho).$
-As the contributions coming from the external potentials can be computed directly with a couple of one-body object $(\rho,n)$, we will now introduce a universal functional of $(\rho,n)$ which contains all necessary information to compute the Gibbs free energy.
-This functional is built using the Levy-Lieb constrained search formalism as in eDFT.
-First, we introduce the intrinsic Gibbs free energy functional of the $N$-body objects
-$f[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]-\text{k}_{\text{B}}T\mathcal{S}[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}].$ (51)
-such that we can then rewrite the Gibbs free energy as
+Since the contributions arising from the external potentials can be directly expressed in terms of the one-body densities $(\rho,n)$, we now introduce a universal functional of $(\rho,n)$ which contains all the information required to compute the free energy.
+This functional is built using the Levy-Lieb constrained search formalism, as in eDFT.
+First, we introduce the intrinsic Helmholtz free energy functional of the $N$-body densities
+$f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ $=E_{{\text{mm}}}[\boldsymbol{p}]+E_{\text{qm}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{{\text{qm}}}^{{\text{mm}}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]-T\mathcal{S}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}].$ (54)
+such that we can then rewrite the Helmholtz free energy as
 $F_{0}=\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}\big(E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]+E_{\text{mm}}^{\text{ext}}[\boldsymbol{p}]+f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\big).$ (55)
@@
 $\min_{\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=\min_{(\rho,n)}\,\,\min_{(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)}F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (56)
-where the notation $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is done over the set of couple of $N$-body objects $(\hat{{\boldsymbol{\rho}}},\boldsymbol{p})$ having the couple of 1-body object given by $(\rho,n)$.
+where the notation $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)$ means that the search is performed over the set of couple of $N_{{\text{qm}}}+N_{{\text{mm}}}$-body densities $(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})$ that yield the 1-body densities $(\rho,n)$.
 Therefore, by remembering that the external parts depends only on the one-body densities (see Eqs.
@@
 $+$ $\min_{(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]\big].$
-Therefore, we can introduce the universal Levy-Lieb functional of the one-body classical and quantum density as
+We have just introduced the universal Levy-Lieb functional of the one-body classical and quantum density,
 $\mathcal{F}[\rho,n]=\min_{(\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p})\rightarrow(\rho,n)}f[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}],$ (58)
 such that one can rewrite the Gibbs free energy as
-$F_{0}=\min_{(\rho,n)}\big{(}(V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big{)},$ (56)
-which achieves to prove that we can obtain $F_{0}$ as a minimization over only one-body densities.
+$F_{0}=\min_{(\rho,n)}\big((V^{\text{ext}}_{{\text{mm}}}|n)+(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}[\rho,n]\big).$ (59)
+This provides a constructive proof of the existence of a variational principle for the Helmholtz Free energy, $F_{0}$, as a functional of one-body densities.
 III.2 Explicit coupling in terms of eDFT and cDFT
 Although Eq.
-(56) allows us to obtain the Gibbs free energy only in terms of the one-body densities $(\rho,n)$, such a functional remains unknown and the explicit connection with well known universal quantum and classical density functionals is yet unclear.
-This paragraph is concerned with a rewriting of $\mathcal{F}[\rho,n]$ such that the functional of MerminMermin (1965) and EvansEvans (1979) explicitly appear, together with a new QM/MM correlation functional.
+(59) expresses the Helmholtz free energy only in terms of the one-body densities $(\rho,n)$, the functional itself remains unknown in practice.
+Thus, the explicit connection with well known universal quantum and classical density functionals is still unclear.
+This paragraph focuses on rewriting of $\mathcal{F}[\rho,n]$ so that the functionals of MerminMermin (1965) and EvansEvans (1979) explicitly appear, along with a new QM/MM correlation functional.
 We begin by recalling the universal purely quantum one-body density functionals of Mermin
-$\mathcal{F}_{\text{qm}}[\rho]=\min_{\hat{{\boldsymbol{\rho}}}\rightarrow\rho}\text{Tr}\{\hat{{\boldsymbol{\rho}}}\big{(}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}})\big{)}\},$ (57)
+[…]
 where $\hat{{\boldsymbol{\rho}}\,}$ is a quantum $N$-body density matrix and $(\hat{{\boldsymbol{\rho}}\,}\rightarrow\rho)$ means that the minimization is performed over $N$-body quantum density matrices $\hat{{\boldsymbol{\rho}}\,}$ leading to a given quantum one-body density $\rho$.
 Similarly, we can define the universal classical one-body functional of Evans
-$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big{(}{\bf W}_{{\text{mm}}}+\text{k}_{\text{B}}T\log(\boldsymbol{p})\big{)}\},$ (58)
+$\mathcal{F}_{\text{mm}}[n]=\min_{\boldsymbol{p}\rightarrow n}\text{Tr}\{\boldsymbol{p}\big({\bf W}_{{\text{mm}}}+T\log(\boldsymbol{p})\big)\},$ (61)
 where $\boldsymbol{p}$ is a classical $N$-body probability distribution and $(\boldsymbol{p}\rightarrow n)$ means that the minimization is performed over $N$-body classical probability distribution $\boldsymbol{p}$ leading to a given classical one-body density $n$.
@@
 $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ $=\int\text{d}{q}\int\text{d}{Q}{W}_{{\text{qm}}}^{{\text{mm}}}({Q},{q})n({Q})\rho({q}),$ (63)
-which will be the QM/MM equivalent of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional defined of course as the complementary functional
+which is the QM/MM analogue of the Hartree-term in eDFT, and where $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ is the QM/MM correlation functional, i.e, the complementary functional
 $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]=\mathcal{F}[\rho,n]-\big(\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big).$ (64)
@@
 We can then rewrite explicitly $F_{0}$ as
-[…]
+$F_{0}=\min_{(\rho,n)}\big\{$ $(v^{\text{ext}}_{{\text{qm}}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]+\mathcal{F}_{\text{mm}}[n]+(V^{\text{ext}}_{{\text{mm}}}|n)$ (65)
 $+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big\},$
@@
 (2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
+We highlight here that, although the outcome of the present derivation is rather natural as it consists of a multi component DFT formalism, the explicit definition of $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq.
+(64) can help in future developments of approximations for $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$.
+For instance, the variational formulation in terms of full-body densities allows the use the adiabatic connection to reformulate the problem, and therefore potentially to import several successful strategies used in eDFT to approximate correlation functionals.
 III.3 Applications to solvation problems
@@
 This is a route that should be explored to improve the quality of the QM/MM model.
-We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $\text{k}_{\text{B}}T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq.
+We now assume that, in the absence of the solvent, the typical electronic gap is large compared to the thermal agitation $T$, allowing us to neglect the contribution of electronic entropy in $\mathcal{F}_{\text{qm}}[\rho]$ (see Eq.
 (60)).
 This is equivalent to considering only the ground state of the electronic quantum system.
-To illustrate the validity of this approximation for systems at room temperature ($T=300\,K$), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$.
+To illustrate the validity of this approximation for systems at room temperature (T=300KT=300\,K), we note that the Boltzmann weight ratio between the ground state and an excited state with a gap of 0.25 eV ($\approx 0.01\,\text{Ha}$) is approximately $10^{-5}$.
 For a gap of 1.63 eV ($\approx 0.06\,\text{Ha}$), which marks the onset of the visible spectrum, this ratio drops dramatically to $10^{-28}$.
@@
 (60)) to recover the usual universal Levy-Lieb electronic density functional, namely
-$\mathcal{F}_{\text{qm}}[\rho]\approx\min_{\Psi\rightarrow\rho}{\langle\Psi|}\hat{\bf t}+\hat{\bf w}_{{\text{qm}}}{|\Psi\rangle}.$ (63)
+[…]
 Regarding solvent effects, we focus on cases where the solvent-solute interaction is not too strong, meaning there is no covalent bonding between the solute and solvent species.
-Under these conditions, the previous $T=0~{}K$ assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
-In previous work about mixed electronic-classical density functional theory to study QM/MM systems, the classical fluid was described in the Grand Canonical ensemble.
-This is because classical DFT is naturally formulated in this ensemble.
-In practice this means that the number of classical particle is now allowed to fluctuate by exchanging with a reservoir of fixed chemical potential $\mu$.
-In this ensemble, Eq.
-(62)
-[…]
-$+\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]+\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]\big{\}}$
-where $\Omega_{0}$ is the grand potential, that is the state function in the grand canonical ensemble, and $\mathcal{F}_{\text{mm}}$ is now the grand canonical universal functional a la Evans.
+Under these conditions, the previous T=0KT=0~K assumption for $\mathcal{F}_{\text{qm}}[\rho]$ remains reasonable, as the electronic structure of the solute is not significantly altered by the presence of the solvent.
+Previous attempts to apply mixed electronic-classical density functional theory to QM/MM systems were carried on in a semi-Grand canonical ensemble, where the number of classical particles is allowed to fluctuate.
+This choice was motivated by the availability of classical functionals, which are naturally formulated in the grand canonical ensemble.
+In the semi-Grand canonical ensemble, the state function is the grand potential, $\Omega_{0}$, for which the following variational principle, analogous to Eq.
+(65), stands
+$\Omega_{0}=$ $\min_{(\rho,n)}\big\{\mathcal{F}_{\text{qm}}[\rho]+\Omega_{\text{mm}}[n^{\text{GC}}]+\delta\Omega_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]$ (67)
+$+\varepsilon_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]+\left(v^{\text{ext}}_{{\text{qm}}}\mid\rho\right)+\left(V^{\text{ext}}_{{\text{mm}}}\mid n^{\text{GC}}\right)\big\}.$
 In Eq.
-(64) one recognizes the usual quantum Born-Oppenheimer part for the electron (i.e.
-$(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$), together with the purely classical part of the solvent (i.e.
-$\mathcal{F}_{\text{mm}}[n]+(v_{nucl}^{{\text{mm}}}-\mu|n)$), and $\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for the mean-field electron-solvent interaction, while $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ accounts for solvation correlation effects in the QMMM system.
-In practice, Eq.
-(64) requires some approximations.
+(67), $n^{\text{GC}}$ is the one-body density of classical particles, $\Omega_{\text{mm}}[n^{\text{GC}}]$ is the grand canonical universal functional a la Evans, $V^{\text{ext}}_{{\text{mm}}}$ is the external classical potential, $(v_{\text{ne}}|\rho)+\mathcal{F}_{\text{qm}}[\rho]$) is the usual quantum Born-Oppenheimer electronic functional and $\delta\Omega_{{\text{qm}}}^{\text{mm}}[\rho,n^{\text{GC}}]$ is the unknown QM/MM correlation term.
+Eq.
+(67) can be obtained following the same route that was used to derive Eq.
+(65).
+Its explicit derivation is in available in S.I.
+III.
+Minimizing Eq.
+(67) in practice require some approximations.
 First, the choice of the functionals describing the quantum system and the liquid considerably influences the quality of the predictions.
-However, determining the appropriate electronic functional to describe a given solute and the suitable functional for a classical fluid are topics that themselves deserve a thorough review.
-Therefore, we let this discussion aside to focus on the QM/MM coupling.
+However, determining the appropriate electronic functional for a given solute and the suitable functional for a classical fluid are complex topics that themselves deserve a thorough review.
+Therefore, we let this discussion aside to focus on the QM/MM coupling and the several strategies adopted in the litterature.
 In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al.
@@
 However, we do not resort to point charges representation to compute the electrostatic energy but directly use the electronic density
-[…]
+$\varepsilon_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ $=\int d{q}\int dr\frac{\rho(q)n_{C}(r)}{|r-{q}|}$ (68)
 where $n_{C}$ is the solvent charge density.
-We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.62.
+We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.
+(65).
 The results obtained in these worksPetrosyan, Rigos, and Arias (2005); Petrosyan et al.
@@
 An important result is that the exact Helmholtz free energy of a QM/MM system can be obtained though a minimization of functionals over objects of reduced dimensions, namely the one-body classical and quantum density.
+The assumptions of these derivations is a QM/MM partitioning of the system, together with models for both the MM part and the QM/MM interaction.
+We therefore do not discuss the QM/MM crossover nor how to go beyond the classical models.
 The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec.
@@
 Supplementary materials
-The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble and II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators.
+The supplementary materials document contains additional information concerning: I) a recall of the variational formulation of statistical quantum mechanics in the canonical ensemble, II) an alternative derivation of the adiabatic QM-MM equilibrium density which does not rely on the Liouville equations or Poisson bracket operators, and III) an extension of the QM/MM functional to the semi-grand-canonical ensemble where the number of classical particles is allowed to vary.
 Appendix A: Properties of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
 We wish to show that i) the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ of Eq.
-(32) is bounded from below by the Helmholtz free energy $F_{0}=-\text{k}_{\text{B}}T\log(Z)$ with $Z$ defined in Eq.
+(35) is bounded from below by the Helmholtz free energy $F_{0}=-T\log(Z)$ with $Z$ defined in Eq.
 (17), and ii) that the only minimizer of $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is $\mathcal{P}_{\text{eq}}$ of in Eq.
@@
 such that the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ can be rewritten as an average value on the measure $\mu({\mathbf{X}})$
-$F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]=$ $\text{Tr}\big{\{}\hat{{\boldsymbol{\rho}}}\boldsymbol{p}\big{(}\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+\text{k}_{\text{B}}T\log(\hat{{\boldsymbol{\rho}}}\boldsymbol{p})\big{)}\}$ (A.12)
-$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big{(}\epsilon({\bf{Q}},i)+\text{k}_{\text{B}}T\log(\mu({\bf{Q}},i))$
-$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big{(}\epsilon({\mathbf{X}})+\text{k}_{\text{B}}T\log(\mu({\mathbf{X}}))$
-$=$ $-\text{k}_{\text{B}}T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
-$=$ $-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
+$F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]=$ $\text{Tr}\big\{\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p}\big(\hat{{\bf V}}^{\text{tot}}_{\text{mm}}+\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}+T\log(\hat{{\boldsymbol{\rho}}\,}\boldsymbol{p})\big)\}$ (A.12)
+$=$ $\int\text{d}{\bf{Q}}\sum_{i}\mu({\bf{Q}},i)\big(\epsilon({\bf{Q}},i)+T\log(\mu({\bf{Q}},i))$
+$=$ $\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\big(\epsilon({\mathbf{X}})+T\log(\mu({\mathbf{X}}))$
+$=$ $-T\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\log(\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})})$
+$=$ $-T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu}$
 $\equiv$ $F[\mu].$
-According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}},\boldsymbol{p}]$ at $\mathcal{P}_{\text{eq}}$ leads straightforwardly to
-$F[\mu_{\text{eq}}]=$ $-\text{k}_{\text{B}}T\log(Z),$ (A.13)
-which is of course the correct Helmholtz free energy. Then two questions remain: i) is $-\text{k}_{\text{B}}T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
+According to Eq. (A.12), the evaluation $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ at $\hat{\mathcal{P}}_{\text{eq}}$ leads straightforwardly to
+$F[\mu_{\text{eq}}]=$ $-T\log(Z),$ (A.13)
+which is of course the correct Helmholtz free energy. Then two questions remain: i) is $T\log(Z)$ a global minima and ii) is $\hat{\mathcal{P}_{\text{eq}}}$ the only minimizer ?
 .2 Use of Jensen inequality to prove variational character
-As the function $-\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer.
+As the function $\log(x)$ is strictly convex, we will now use Jensen’s inequality from convex analysis and probability measures to prove that $F[\mu]$ is bounded from below by the Helmholtz free energy and that $\mu_{\text{eq}}$ is the only minimizer.
 The Jensen’s inequality can be stated as follows.
@@
 $\varphi\bigg(\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})g({\mathbf{X}})\bigg)\leq\int\text{d}{\mathbf{X}}\mu({\mathbf{X}})\varphi\big(g({\mathbf{X}})\big).$ (A.15)
-As $-\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
+As $\log(x)$ is convex, the Jensen’s equality can be applied to $F[\mu]$ of Eq. (A.12) by setting
 $g({\mathbf{X}})=\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})},$ (A.16)
-$\varphi(x)=-\text{k}_{\text{B}}T\log(x).$
+$\varphi(x)=-T\log(x).$
 Therefore the inequality of Eq. (A.14) becomes
-$-\text{k}_{\text{B}}T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-\text{k}_{\text{B}}T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
+$T\log\langle\frac{e^{-\beta\epsilon}}{\mu}\rangle_{\mu}\leq-T\langle\log(\frac{e^{-\beta\epsilon}}{\mu})\rangle_{\mu},$ (A.17)
 and as
@@
 one can rewrite Eq. (A.17) as
-$-\text{k}_{\text{B}}T\log(Z)\leq F[\mu],$ (A.19)
+$-T\log(Z)\leq F[\mu],$ (A.19)
 which guarantees the variational character.
-Eventually, the strict convexity of the $-\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into
+Eventually, the strict convexity of the $\log(x)$ function implies that the Jensen’s inequality becomes an equality if and only if $g({\mathbf{X}})=cst$, which translates into
 $g({\mathbf{X}})=cst\Leftrightarrow\frac{e^{-\beta\epsilon({\mathbf{X}})}}{\mu({\mathbf{X}})}=cst\Leftrightarrow\mu({\mathbf{X}})\propto e^{-\beta\epsilon({\mathbf{X}})},$ (A.20)

### v6

@@
 Combining classical density functional theory (cDFT) with quantum mechanics (QM) methods offers a computationally efficient alternative to traditional QM/molecular mechanics (MM) approaches for modeling mixed quantum-classical systems at finite temperatures.
-However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region
+However, both QM/MM and QM/cDFT rely on somewhat ambiguous approximations , the two major ones being: i) the definition of the QM and MM regions as well as the description of their coupling, and ii) the choice of the methods and levels of approximation made to describe each region.
 This paper addresses the second point and develop an exact theoretical framework that allows us to clarify the approximations involved in the QM/cDFT formulation.
 We therefore establish a comprehensive density functional theory (DFT) framework for mixed quantum-classical systems within the canonical ensemble.
-We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles.
+We start by recalling the expression of the adiabatic equilibrium density matrix for a mixed system made of $N_{{\text{qm}}}$ quantum and $N_{{\text{mm}}}$ classical particles, which can be related to a partial Wigner transformation.
 Then, we propose a variational formulation of the Helmholtz free energy in terms of the full , non-equilibrium, QM/MM density matrix.
@@
 from the solutions of the Born Oppenheimer Hamiltonian.
-This forms the basis of ab-initio molecular dynamics Iftimie, Minary, and Tuckerman (2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
+This forms the basis of ab-initio molecular dynamics Iftimie et al.
+(2005); Silvestrelli and Parrinello (1999) which recovers the classical nuclear entropic contributions through either classical Monte Carlo algorithms or classical trajectories under the ergotic hypothesis.
 Nevertheless, it is important to keep in mind that nuclear quantum effects play a significant role for light nuclei and can substantially alter the rates of certain chemical processes, such as proton transfer reactionsBała et al.
-(1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin, Robert, and Savéant (2010); Hammes-Schiffer and Stuchebrukhov (2010).
+(1996); Hanna and Kapral (2005); Kim and Hammes-Schiffer (2006); Costentin et al.
+(2010); Hammes-Schiffer and Stuchebrukhov (2010).
 Moreover, such proton transfers can also be strongly influenced by the solvation effects, specially when polar solvents are usedKapral (2015).
@@
 (2018).
-To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland, Lyne, and Karplus (2000); Senn and Thiel (2009a); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM).
+To alleviate ab-initio molecular dynamics computational burden, QM/MMWarshel and Levitt (1976); Mulholland et al.
+(2000); Senn and Thiel (2009); Groenhof (2013) methods were developed, where only a small, critical portion of the system is treated at the QM level, while the remainder is handled at the classical level (MM).
 The interaction between the classical particles are modeled using parametrized force fields and their dynamics obey Newton’s law of motion.
@@
 (2021); Wu et al.
-(2023); Böselt, Thürlemann, and Riniker (2021); Ho et al.
+(2023); Böselt et al.
+(2021); Ho et al.
 (2025) involving both physically-motivated and machine-learning oriented approaches.
@@
 Its computational efficiency makes cDFT appealing to model solvent effect at the classical level while retaining a quantum description of the soluteTang et al.
-(2020); Petrosyan, Rigos, and Arias (2005); Petrosyan et al.
-(2007); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
+(2020); Petrosyan et al.
+(2005, 2007); Jeanmairet et al.
+(2020); Labat et al.
+(2024).
 In this approach, the total Hamiltonian is partitioned into well defined QM and MM Hamiltonians.
 However, just as in molecular dynamics based QM/MM approaches, the coupling interaction between the QM and MM regions lacks of a unique practical definition.
-The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009b); Giovannini, Lafiosca, and Cappelli (2017); Clemente, Capece, and Martí (2023).
+The coupling Hamiltonian is usually described as the sum of an electrostatic interaction combined with a non-bonded term which is often modeled using Lennard-Jones force field to account for repulsion and dispersion effectsSenn and Thiel (2009); Giovannini et al.
+(2017); Clemente et al.
+(2023).
 The Mixed cDFT-eDFT approach advantageously replaces the exhaustive phase-space sampling required in QM/MM molecular dynamics with a self-consistent optimization of both classical and quantum functionals.
 While the idea of merging cDFT and quantum methods is completely natural its theoretical justification is not immediately straightforward.
-Petrosyan and co-workersPetrosyan, Rigos, and Arias (2005); Petrosyan et al.
-(2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem.
+Petrosyan and co-workersPetrosyan et al.
+(2005, 2007) introduced the joint density functional theory which consists in a DFT formulation of the QM/MM problem.
 The authors obtain a multi component electron/nuclei/solvent DFT formulation, which is justified by the use of a Levy-Lieb constrained search procedure allowing to express the universal functional as depending on the electron density and the nuclear density rather than the full density matrix describing the electrons and nuclei.
@@
 Their working functional is instead defined as the sum of the Quantum Kohn-Sham functional describing the electrons of the solute, the classical functional describing the solvent and a coupling functional defined as the difference between the exact functional and the sum of the two former functionals.
-In other workJeanmairet, Levesque, and Borgis (2020); Tang et al.
+In other workJeanmairet et al.
+(2020); Tang et al.
 (2020), a similar form of the functional was invoked without further theoretical justification.
@@
 Under these assumptions, the primary objective of this paper is to derive, in a straightforward manner, the one-body density functional theory (DFT) formulation for a QM/MM system.
-Therefore, The key steps of this derivation are as follows: i) We begin by defining the general Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian (see Sec.
-II.1).
+Therefore, The key steps of this derivation are as follows: i) We begin by a qualitative discussion in Sec.
+II.1 sketching how the mathematical objects of QM/MM systems can be obtained by Wigner transformations, and then define properly in Sec.
+II.2 the general form of the Hamiltonian of the system, considering a broadly applicable QM/MM Hamiltonian.
 ii) We then express the exact solution of the problem using full-system quantities (i.e., those describing all particles).
 This includes the Helmholtz free energy of the QM/MM system, along with the corresponding equilibrium density matrix and partition function (see Sec.
-II.2).
+II.3).
 iii) Next, we reformulate the problem as the minimization of a functional over generic full-system density matrices.
 In our case, this means expressing the Helmholtz free energy as a functional of QM/MM density matrices that describe all particles in the system (see Sec.
-II.3.1).
-iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.3.2).
+II.4.1).
+iv) We then decompose this functional into intrinsic (arising from kinetic energy, inter-particle interactions and entropic contributions) and external (arising from external potentials) contributions (see Sec II.4.2).
 v) Under certain assumptions about the form of the external potentials, we show how the external part of the functional can be written as a linear functional of one-body densities.
@@
 al.
-in the context of QM/MM dynamicsNielsen, Kapral, and Ciccotti (2001), and an extension to the semi grand canonical ensemble of the present formalism.
+in the context of QM/MM dynamicsNielsen et al.
+(2001), and an extension to the semi grand canonical ensemble of the present formalism.
 As this work is intended for both the QM and MM communities, which often differ in terminology and conceptual frameworks, we highlight certain elements throughout the text to aid accessibility.
@@
 This variational principle, although hardly usable in practice due to the high-dimensionality of the objects it manipulates, is an important milestone to properly obtain the one-body formulation, which is the aim of this paper.
-We define in Sec.
-II.1 the working Hamiltonian that describes the QM/MM system.
+As the formalism uses mixed QM/MM objects that might seem rather odd in their mathematical design, we begin this derivation by providing in Sec.II.1 a qualitative discussion on how one can obtain these QM/MM objects thanks to the concept of Wigner transform.
+We then define in Sec.
+II.2 the working Hamiltonian that describes the QM/MM system.
 Again, we do not discuss the validity of the partitioning or the quality of the coupling term, we assume that these preliminary considerations have already been addressed.
 The equilibrium density of a mixed QM/MM system is then given in Sec.
-II.2, and we briefly discuss the physical meaning of that result.
+II.3, and we briefly discuss the physical meaning of that result.
 We eventually show in Sec.
-II.3 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
-II.1 The Hamiltonian of the quantum-classical system
+II.4 that the Helmholtz free energy can be obtained through a variational principle over mixed QM/MM density matrices.
+II.1 Qualitative discussion: dealing with both QM and MM particles thanks to the Wigner transform
+Properly defining objects for a QM/MM system can appear as a priori relatively odd as the mathematical spaces of the two theories are very different.
+We provide here a qualitative discussion (which is fully expanded in the supplementary materials) on how one can obtain naturally QM/MM objects with the help of the so-called Wigner-Weyl transformation.
+Observables in classical statistical mechanics are obtained as integrals over the $(Q,P)$ classical phase space,
+$\langle O\rangle_{f}=\int dQdPf(Q,P)O(Q,P),$ (1)
+where $f(Q,P)$ is a classical probability distribution and $O(Q,P)$ is a classical observable, while the QM theory uses density matrices $\hat{\rho}$ and potentially non local operators $\hat{O}$ acting in $\mathbb{R}\times\mathbb{R}$, namely
+$\langle\hat{H}\rangle_{\hat{\rho}}=\int dxdx^{\prime}\hat{\rho}(x,x^{\prime})\hat{O}(x^{\prime},x),$ (2)
+where $\hat{\rho}(x,x^{\prime})$ and $\hat{O}(x^{\prime},x)$ are the real-space representation of the density matrix $\hat{\rho}$ and $\hat{O}$ operators, respectively.
+Nevertheless, one can notice that the $(Q,P)$ classical phase space is isomorph to the $(x,x^{\prime})$ representations of the QM operators.
+The Wigner-Weyl transformWigner (1932) exploits such a similarity and allows then for a phase space representation of QM density matrices and operators (see Ref.
+Case, 2008 for a pedagogical introduction).
+More precisely, for any QM operator $\hat{O}$ one can define its Wigner transformation labelled $\tilde{O}_{W}$, which is now a function of the $(Q,P)$ variables.
+One can also define the Wigner transformed density matrix $\tilde{\rho}_{W}(Q,P)$ associated to $\hat{\rho}$, which allows then to write the QM observable of Eq.
+(2) as a classical phase space integral, namely
+$\langle\hat{O}\rangle_{\hat{\rho}}=\int dQdP\tilde{\rho}_{W}(Q,P)\tilde{O}_{W}(Q,P).$ (3)
+An important property of this framework is that QM operators expressed as ”simple” functions of the QM operators $\hat{Q}$ and $\hat{P}$ (i.e.
+functions with no products of $\hat{P}$ and $\hat{Q}$) are transformed as the corresponding purely classical function.
+An important special case of this property is the Hamiltonian operator whose Wigner transformation is precisely the classical Hamiltonian, i.e.
+$\tilde{H}_{W}(Q,P)=\frac{P^{2}}{2M}+V(Q)\equiv H(Q,P).$ (4)
+Therefore, the expectation value of the QM Hamiltonian over a QM density matrix is obtained as
+$\langle\hat{H}\rangle_{\hat{\rho}}=\int dXdP\tilde{\rho}_{W}(Q,P)H(Q,P),$ (5)
+which is very similar to the definition of the statistical classical mechanics.
+Nevertheless the quantum nature of the problem remains encoded in the fact that the object $\tilde{\rho}_{W}(Q,P)$ is potentially complex-valued in opposition to classical probability distributions.
+[…]
+In the case now where two types of particles are present in the system with a clear difference of masses, one can perform a Wigner transformation only on the heavy particles and one obtains a power series in terms of the ratio of the masses $\alpha\ll 1$.
+As shown originally by Nielsen et.
+al.Nielsen et al.
+(2001), the zeroth-order term in $\alpha$ yields to a classical description of the heavy particles, while maintaining the full QM description of the light particles.
+The obtained semi-Wigner transformed density matrix corresponds then to a QM/MM description of the system: the heavy particles are described by the classical variables $(Q,P)$ while the light particles are described by a non local operator.
+An important aspect of these results is that the QM operators depend parametrically on the $(Q,P)$ classical variables.
+We will use this important result as the starting of our derivation (see Sec.
+II.3) which will allow us to understand how generic QM/MM density matrices can be built in order to design our variational principle.
+II.2 The Hamiltonian of the quantum-classical system
 We consider a QM/MM system, where the partition and the choice of the model to describe the MM part and the coupling term have already been made
@@
 $\hat{\bf H}_{\text{tot}}({\bf{Q}},{\bf P},{\bf{q}})=\hat{\bf h}_{{\text{qm}}}({\bf{q}})+{\bf H}_{{\text{mm}}}({\bf{Q}},{\bf P})+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}},{\bf{q}}).$ (6)
+Notice that the form of the QM/MM Hamiltonian as written in Eq.
+(6) can actually be obtained as the partial Wigner transformation of a purely QM Hamiltonian, and where the Wigner transformation is done only on the $N_{{\text{mm}}}$ particles.
 In the rest of the paper, explicit dependency on the QM and MM variables will be dropped for concision when unambiguous.
@@
 The function ${W}_{{\text{qm}}}^{{\text{mm}}}({q},{Q})$ represents the effective interaction between an effective particle (such as a nuclei, an atom or a molecule) at position Q{Q} and a fundamental particle (such as an electron) at position q{q}.
-II.2 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
+II.3 Mixed quantum-classical equilibrium density, partition function and Helmholtz free energy
 We start by defining the exact equilibrium density matrix (EDM) of the QM/MM system in the canonical ensemble.
@@
 (14) corresponds to the real-space representation of the EDM, originally obtained by Nielsen et.
-al.Nielsen, Kapral, and Ciccotti (2001) in the adiabatic basis, i.e.
+al.Nielsen et al.
+(2001) in the adiabatic basis, i.e.
 in the basis of eigenvectors of $\hat{\bf h}_{{\text{qm}}}+\hat{\bf W}_{{\text{qm}}}^{{\text{mm}}}({\bf{Q}})$.
 In this basis, the density matrix is diagonal , as found by Nielsen et.
-al.Nielsen, Kapral, and Ciccotti (2001).
+al.Nielsen et al.
+(2001).
 In the case of quantum electrons and classical nuclei, the adiabatic states are the eigenstates of the Born-Oppenheimer Hamiltonian.
@@
 This will allow to formulate a variational formulation of the Helmholtz free energy for a QM/MM system in Sec.
-II.3.
+II.4.
 Last but not least, the decomposition of the equilibrium density of Eq.
@@
 Therefore, the two systems are correlated.
-II.3 Variational formulation of the Helmholtz free energy for a QM/MM system
+II.4 Variational formulation of the Helmholtz free energy for a QM/MM system
 Both classical and quantum DFT can be formulated through the Levy-Lieb constrained-search formalism.
@@
 To the best of our knowledge, this was never formalized before.
-II.3.1 Variational formulation
+II.4.1 Variational formulation
 As we know that the equilibrium density for a mixed QM/MM system is expressed as a special case of product of a quantum density matrix by a classical probability distribution, we can now reformulate the Helmholtz free energy $F_{0}$ as a variational problem over QM and MM objects, such that the equilibrium density of Eq.
@@
 III.3.
-II.3.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
+II.4.2 Rewriting of the functional $F[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$
 Because the function $\boldsymbol{p}$ does not explicitly depend on the quantum variables and because we chose the normalization conditions of Eq. (37), we can naturally decompose the entropy as
@@
 $=\int\text{d}{\bf{Q}}\boldsymbol{p}({\bf{Q}})$
-$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+{w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
+$\quad\int\text{d}{\bf{q}}\text{d}{\bf{q}}^{\prime}(\bf{t}({\bf{q}},{\bf{q}}^{\prime})+\delta({\bf{q}}-{\bf{q}}^{\prime}){w}_{{\text{qm}}}({\bf{q}},{\bf{q}}^{\prime})){\boldsymbol{\rho}}({\bf{q}},{\bf{q}}^{\prime},{\bf{Q}}),$
 $E_{\text{qm}}^{\text{ext}}[\hat{{\boldsymbol{\rho}}\,},\boldsymbol{p}]$ is the functional accounting for the interaction of the quantum particles with the external potential
@@
 where we can explicitly recognize the purely quantum and purely classical free-energy functionals, together with the QM/MM coupling expressed as the sum of the mean-field and correlation terms.
-If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan, Rigos, and Arias (2005); Petrosyan et al.
-(2007); Tang et al.
-(2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
+If $\delta\mathcal{F}_{{\text{qm}}}$ is ignored, we recover the mean-field coupling between the eDFT and cDFT functionals that is generally used in the literaturePetrosyan et al.
+(2005, 2007); Tang et al.
+(2020); Jeanmairet et al.
+(2020); Labat et al.
+(2024).
 We highlight here that, although the outcome of the present derivation is rather natural as it consists of a multi component DFT formalism, the explicit definition of $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}[\rho,n]$ in Eq.
@@
 Therefore, we let this discussion aside to focus on the QM/MM coupling and the several strategies adopted in the litterature.
-In Petrosyan et al’s joint DFT Petrosyan, Rigos, and Arias (2005); Petrosyan et al.
-(2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential.
+In Petrosyan et al’s joint DFT Petrosyan et al.
+(2005, 2007), the interaction of the solvent with both the solute nuclei and electrons is dealt as a whole and the coupling is described by the use of pseudo potential.
 Thus Eq.
@@
 Moreover, since both density are not optimized self-consistently the interaction term $\varepsilon_{{\text{qm}}}^{{\text{mm}}}$ only influences the classical optimization and not the quantum one.
-In our previous worksJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq.
+In our previous worksJeanmairet et al.
+(2020); Labat et al.
+(2024), we also ignored the correlation functional $\delta\mathcal{F}_{{\text{qm}}}^{{\text{mm}}}$ in Eq.
 (70) and limit ourselves to the mean-field term.
@@
 where $n_{C}$ is the solvent charge density.
-We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat, Giner, and Jeanmairet (2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.
+We also proposed to couple wave function theory-based methods, such as Hartree-Fock and selected configuration interaction with cDFTLabat et al.
+(2024) to approximate the universal Levy-Lieb functional $\mathcal{F}_{\text{qm}}[\rho]$ of Eq.
 (70).
-The results obtained in these worksPetrosyan, Rigos, and Arias (2005); Petrosyan et al.
-(2007); Tang et al.
-(2020); Jeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
+The results obtained in these worksPetrosyan et al.
+(2005, 2007); Tang et al.
+(2020); Jeanmairet et al.
+(2020); Labat et al.
+(2024) clearly indicate that, for ground state properties such as dipoles for instance, the mean-field QM/MM approximation gives already a decent improvement with respect to in vaccuum pure QM calculations.
 IV Conclusion
@@
 The starting point of these derivations is the proposal of the QM/MM equivalent to the quantum ground state variational principle or the classical Gibbs measure variational principle (see Sec.
-II.3).
+II.4).
 This QM/MM functional is formalized in terms of a $N_{{\text{mm}}}$ classical probability density and an $N_{{\text{qm}}}$ quantum density matrix which also parametrically depends on the $N_{{\text{mm}}}$ classical variables (see Sec.
@@
 (19) and (22)) which can be obtained from first principle through a semi classical expansion of a partial Wigner transformation (see Ref.
-Nielsen, Kapral, and Ciccotti, 2001 and the section II of the supplementary information for a detailed derivation).
+Nielsen et al., 2001 and the section II of the supplementary information for a detailed derivation).
 Starting from this QM/MM variational principle written in terms of $N$-body objects and thanks to constrained search formalisms similar to that proposed by Levy and Lieb in eDFT, we establish a universal QM/MM functional depending only on the one-body classical and quantum objects (see Sec.
@@
 We then propose a practical application of this theory in the case of solvation problems, and propose a mean-field treatment.
-Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet, Levesque, and Borgis (2020); Labat, Giner, and Jeanmairet (2024).
+Thanks to these derivations, it becomes clear that the explicit correlation between the classical and quantum electronic systems, were implicitly neglected in our previously reported QM/cDFT calculationsJeanmairet et al.
+(2020); Labat et al.
+(2024).
 Supplementary materials